### v2

# sPhinX: Sample Efficient Multilingual Instruction Fine-Tuning Through N-shot Guided Prompting

Abstract
Despite the remarkable success of LLMs in English, there is a significant gap in performance in non-English languages. In order to address this, we introduce a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX, which is created by selectively translating instruction response pairs from English into 50 languages. We test the effectiveness of sPhinX by using it to fine-tune two state-of-the-art models, Phi-3-small and Mistral-7B and then evaluating them across a comprehensive suite of multilingual benchmarks that test reasoning, question answering, and reading comprehension. Our results show that Phi-3-small and Mistral-7B fine-tuned with sPhinX perform better on an average by 4.2%pt and 5%pt respectively as compared to the baselines. We also devise a strategy to incorporate N-shot examples in each fine-tuning sample which further boosts the performance of these models by 3%pt and 10%pt respectively. sPhinX outperforms other multilingual instruction tuning datasets on the same benchmarks along with being sample efficient and diverse, thereby reducing dataset creation costs. Additionally, instruction tuning with sPhinX does not lead to regression on most standard LLM benchmarks.
1 Introduction
Large Language Models (LLMs) have been shown to perform several tasks exceptionally well in English, however, performance in some non-English languages still lags behind Ahuja et al. (2023); Asai et al. (2023). Further, the gap between the performance of Large Language Models (LLMs) and Small Language Models (SLMs) is more pronounced Ahuja et al. (2024) on
non-English languages. Cui et al. (2023) and Balachandran (2023) follow the approach of training the models further on datasets for specific languages, however this can lead to catastrophic forgetting and hurt the performance on English Zhao et al. (2024). Several techniques have been proposed to bridge this gap, such as incorporating better pre-training data in various languages and enhancing base tokenizers Xu et al. (2024); Dagan et al. (2024). However, most of these changes need to be implemented at the pre-training stage which demands extensive data and computational resources, making it practically unfeasible in many scenarios Brown et al. (2020). Consequently, the most well-studied technique involves fine-tuning models for specific languages and tasks. Instruction fine-tuning (IFT) has become a popular technique for enhancing the performance of language models in specific languages. This method combines the benefits of both the pretrain–fine-tune and prompting paradigms Wei et al. (2021).
Sample diversity is crucial for instruction tuning in multilingual datasets. Many recent datasets are created by translating English content into other languages or using self-instruct techniques from seed prompts Li et al. (2023); Taori et al. (2023), both of which can limit diversity. Machine translation can lead to loss of semantic information Baroni and Bernardini (2006), while self-instruct methods may produce repetitive and homogeneous samples Wang et al. (2022). This underscores the importance of having a dataset with diverse set of samples.
Aggarwal et al. (2024) evaluate several models fine-tuned using Parameter Efficient fine-tuning and find that there is a gain in multilingual performance across several SLMs for many low-resource languages, with some high-resource languages performing worse after fine-tuning. However, the performance gains often do not match the performance of larger models, such as GPT-4 and Gemini and can be inconsistent across languages. Hence, there is a need to study instruction tuning for better multilingual performance in SLMs.
In this paper, we present a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX.  It comprises 1.8M instruction-response pairs in 51 languages, derived by selectively translating the Orca instruction tuning dataset Mukherjee et al. (2023) using GPT-4 Achiam et al. (2023). We assess the effectiveness of sPhinX by fine-tuning two models — Phi-3-small and Mistral-7B — across a range of evaluation benchmarks that test various language model capabilities like reasoning, question answering, and reading comprehension. We compare models fine-tuned on sPhinX with those using other synthetic multilingual instruction tuning datasets like Aya Üstün et al. (2024), Multilingual Alpaca Taori et al. (2023), and Bactrian Li et al. (2023) and observe significant performance gains.
The contributions of this paper are as follows:
  • We introduce a novel approach to generate synthetic data for multilingual instruction tuning by selectively translating the Orca dataset with the assistance of GPT-4 (§3.1)
  • We devise LAnguage-Specific N-shot Guided Instruction fine-tuning (LANG) strategy for enhancing the multilingual capabilities of LLMs (§4.2)
  • We conduct an extensive number of experiments on different multilingual instruction tuning datasets to test generalizability in multilingual settings (§5)
  • We also conduct an in-depth analysis where we find sPhinX to be more sample efficient and diverse across languages (§3.3)
2 Related Work
2.1 Multilingual LLMs
Recently, there has been there has been interest in developing and improving SLMs such as LLaMA Touvron et al. (2023), Mistral Jiang et al. (2023), Phi3 Abdin et al. (2024), and Gemma Team et al. (2024). In a relatively brief period, researchers have developed numerous monolingual and multilingual versions of these foundational models, either by pre-training from scratch on specific languages or groups of languages, or by further fine-tuning on them Nguyen et al. (2023); Gala et al. (2024); Balachandran (2023); Cui et al. (2023); Zhang et al. (2023); Qin et al. (2024). Prior work has shown that smaller models show promising results in high-resource languages but perform worse on low-resource ones Ahuja et al. (2024) compared to larger models. To address this, researchers have proposed using multilingual corpora for pre-training, expanding the vocabulary Wang et al. (2019) and continual training Liu et al. (2023), though this increases training costs significantly. fine-tuning for specific tasks, such as translation, has been explored but often struggles with generalization  Mishra et al. (2021). In contrast, instruction tuning has shown to improve generalization to unseen tasks, enhancing the models’ ability to understand and respond to human instructions  Mishra et al. (2021); Ouyang et al. (2022). This makes the models more versatile and capable of handling a wide variety of tasks.
2.2 Multilingual Instruction fine-tuning
Early studies focused on fine-tuning pre-trained models on a variety of languages through data augmentation for a single task Hu et al. (2020); Longpre et al. (2021); Asai et al. (2022). Currently, the approach has shifted on fine-tuning these models using a wide-variety of tasks Longpre et al. (2023); Ouyang et al. (2022). Models such as BLOOMZ Muennighoff et al. (2022) and mT0 Muennighoff et al. (2022) makes significant strides in improving the multilingual performance of decoder-based models Ahuja et al. (2023). There have been multiple multilingual instruction datasets and models proposed such as Bactrian Li et al. (2023), AYA Üstün et al. (2024), polyLM Wei et al. (2023b) after BLOOMZ and mT0. However, these models still do not perform as well as English on other languages, with the gap being particularly large for low-resource languages and languages written in scripts other than the Latin script  Ruder et al. (2021); Ahuja et al. (2023); Asai et al. (2023); Ahuja et al. (2024). In this work, we aim to narrow the performance gap by introducing a strategy for creating datasets for multilingual instruction tuning and recipes for fine-tuning, which we will discuss in the following sections.
Figure 1: The above figure describes translating using a Translation API vs Selective Translation
2.3 Multilingual Synthetic Data Generation
Most instruction-tuning datasets across multiple languages typically focus on general tasks rather than specific reasoning capabilities. While datasets like Orca Mukherjee et al. (2023) and Orca 2 Mitra et al. (2023) exist in English, they highlight a prevalent issue: current methods often prioritize style imitation over leveraging the reasoning abilities found in large foundation models (LFMs). The Orca dataset addresses this by imitating rich signals from GPT-4, including explanation traces and step-by-step thought processes Wei et al. (2023a), guided by assistance from ChatGPT. In order to create multilingual datasets, researchers commonly use translation APIs or LLMs to translate English-specific datasets into target languages. For example, the Bactrian dataset Li et al. (2023) translates Alpaca and Dolly instructions into 52 languages using the Google Translator API and generates outputs with GPT-3.5 turbo. However, these translated datasets often struggle to encode semantic information effectively Baroni and Bernardini (2006). Our dataset approach aims to tackle these challenges by selectively translating only the essential portions of multilingual inputs. This strategy not only preserves semantic information but also accommodates diverse linguistic contexts, thereby enhancing the overall quality and applicability of instruction-tuning datasets across languages.
3 sPhinX Dataset
In this section, we describe our dataset construction methodology (§3.1), dataset filtering and cleaning pipelines (§3.2), and exploratory analysis (§3.3) for determining diversity of our data in comparison to other instruction tuning datasets.
Figure 2: Some examples of input queries and its counterpart existing in the hindi version of the Multialpaca dataset and if it was generated using the Selective Translation strategy
3.1 Dataset Construction
Mukherjee et al. (2023) illustrate how smaller models can replicate the reasoning processes of LLMs and learn from detailed signals like as explanation traces, step-by-step thought processes, and other intricate instructions obtained from data annotated by ChatGPT and GPT-4111GPT-4 inference hyper-parameters in Azure OpenAI interface set as: temperature=0.0, top_p=1, frequency_penalty=0, presence_penalty=0, stop=[”¡—im_end—¿”, ”¡—im_start|¿”]. Their dataset consists of 5M ChatGPT and 1M GPT-4 generated instruction-response pairs. Inspired by this work, we utilized 1M GPT-4 generated instruction-response pairs of their dataset and construct our dataset along similar lines by selectively translating them into 50 different languages with the help of GPT-4. We categorize them into three groups: high-resource, mid-resource, and low-resource languages as outlined in Table 11. For high resource languages, we randomly sample 100k instruction-response pairs from the Orca 1M dataset and generate the responses from GPT-4 using Selective Translation as shown in Figure 1. Similarly, we leverage the same strategy for medium and low resource languages by sampling 50k and 25k pairs respectively. Although GPT-4 performs competitively with commercial translation systems (Google Translate & Bing Translate) it still lags behind on medium and low resource languages Jiao et al. (2023); Hendy et al. (2023). Furthermore, as highlighted in Chang et al. (2023), fine-tuning with a large set of samples from medium and low-resource languages might lead to catastrophic forgetting of high-resource languages. Therefore, we deliberately create fewer samples for medium and low-resource languages than for high-resource ones.
A fundamental problem with using an off-the-shelf translation API is the lack of semantic and task awareness, in addition to translationese Baroni and Bernardini (2006)), which can result in poor quality training data. Consider for example the task of Machine Translation as part of the instruction, wherein the language of the source sentence should be retained. However, an off-the-shelf API, without task awareness, would translate it, resulting in an ill defined instruction. To mitigate this issue, we prompt GPT-4 to selectively translate the instructions, so that the tasks are translated into the appropriate language without changing the semantic meaning. Figure 2 illustrates this with concrete examples. The first example demonstrates the aforementioned translation inconsistency issue for an instruction asking for a French equivalent of an English phrase. The second example demonstrates the consequence of direct translations in the MultiAlpaca dataset: wherein the translation of the task input results in the task being ill-defined based on the instructions. As demonstrated, our proposed Selective Translation method is able to keep the semantic information of the task intact while translating the instructions. For the exact prompt, please refer to Figure 6 in the Appendix.
3.2 Dataset Filtering and Cleaning
Input: listOfSentences: list of strings
1 Function dataFilter(listOfSentences):
2       englishWords $\leftarrow$ set of English words from NLTK;
3       foreach sentence in listOfSentences do
4             cleanedSentence $\leftarrow$ replace all punctuations, digits, and single characters with a single space;
5             cleanedSentence $\leftarrow$ replace all sequences of whitespace with a single space;
6             wordCount $\leftarrow$ 0;
7             foreach word in cleanedSentence do
8                   if word.lower() in englishWords then
9                         wordCount $\leftarrow$ wordCount + 1;
10
11
12
13            content $\leftarrow$ wordCount / len(cleanedSentence);
14             if content $>$ 0.90 then
15                   remove sentence from listOfSentences
16
17
Algorithm 1 Data Filtering Algorithm
Figure 3: t-SNE Visualisation of 1000 samples equally distributed in 10 languages
After generating these translations, we analyze how many instruction-assistant pairs were unsatisfactorily translated, such as incomplete or missing translations, by identifying the English content within the instances. We use the NLTK [^2] https://www.nltk.org/ library to identify the fraction of English words in the sample. The NLTK corpus contains approximately 240,000 English words, making it suitable for our use case. We manually analyzed various proportions of English content in samples of each language and found that the samples were satisfactory when the English content was within 90% for all languages. If the English content exceeded 90% in a language that uses a different script, those samples were removed from our dataset as shown in our data filtering algorithm here. Finally, we manually inspected a small sample of the dataset to ensure the quality of the translations, and found them to be of good quality.
After filtering, the final dataset comprised of 1.8 million samples in 51 languages, divided into three subsets: Train, Test, and Few-shot. We partitioned each language’s dataset to ensure the Test and Few-shot sets contained 2000 and 1000 samples respectively, with the Train set consisting the remaining data. This approach guarantees consistent distributions across languages in the Test and Few-shot sets, ensuring equitable representation regardless of training distribution. The train, test, and few-shot sets are in the ratio of 92 : 5.3 : 2.7.
3.3 Data Analysis
A primary issue that we hypothesize (and find preliminary evidence for from basic analysis of prior Multilingual IFT datasets) is the lack of diversity of samples. In order to validate this issue, we use t-Distributed Stochastic Neighbor Embedding (t-SNE) Cai and Ma (2022) to visualize the structure of both prior multilingual IFT datasets as well as sPhinX. We use the model paraphrase-multilingual-MiniLM-L12-v2 from huggingface’s sentence-transformers Reimers and Gurevych (2020), Reimers and Gurevych (2019) library [^3] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 to generate the embeddings. The hyperparameters were set as follows: learning rate at 200, perplexity at 30, and number of iterations (n_iter) at 1000. Concretely, we use 1000 samples equally distributed in 10 languages. Figure 3 demonstrates that sPhinX demonstrates greater diversity compared to prior datasets. Particularly, Bactrian and Multilingual Alpaca appear less diverse, which we hypothesize might be due to their generation method involving translations of the same set of task limited English samples into different languages. By virtue of using a more diverse seed set (Mukherjee et al., 2023), we circumvent this issue by design. Furthermore, one notable differentiating factor of sPhinX is the substantial amount of code-switched data the inherent nature of Selective Translation. This in turn enhances the diversity of samples across multiple languages. Finally, we also observe the Aya dataset to be more diverse than other prior counterparts, though it still shows less diversity across task instructions.
Dataset Average Token
Length/Sample
Aya 2240
Bactrian 2465
Multilingual Alpaca 1620
sPhinX-0s 544
sPhinX 3100
Table 1: Average Token Length in each dataset
4 Experiments
4.1 Setup
Base Models: We use Mistral-7B  [^4] We specifically use the v1.0 base model from here. Jiang et al. (2023) and Phi-3-small Abdin et al. (2024) base model variants and fine-tune them on different datasets based to improve multilingual performance.
Datasets: Apart from the sPhinX dataset, we use Bactrian Li et al. (2023), MultiAlpaca Wei et al. (2023b) and Aya Singh et al. (2024) instruction datasets for comparative evaluation §3.1.
  • Bactrian Li et al. (2023) is a machine translated dataset of the original alpaca-52k Taori et al. (2023) and dolly-15k Conover et al. (2023) datasets into 52 languages. The instructions for this dataset were translated using a Translation API and then GPT-3.5-Turbo was prompted to generate outputs. We fine-tune our models on the complete dataset consisting of 3.4M instances.
  • MultiAlpaca Wei et al. (2023b) is a self instruct dataset which follows the same approach as the English-only Alpaca by translating the seed instructions to 11 languages and then using GPT-3.5.-Turbo to generate responses. We fine-tune our models on the complete set of the dataset consisting of 500k datapoints.
  • Aya Singh et al. (2024) consists of human-curated prompt-completion pairs in 65 languages called the Aya Dataset. It also consists of an aggregation of 44 monolingual and multilingual templated instruction datasets and 19 translated datasets ranging over 114 languages. The total size of the Aya dataset is around 513M instances. Due to the in-feasibility of fine-tuning our models on the entire dataset, we sampled it down to 2.7M instances, ensuring parity with our sPhinX dataset by selecting equal numbers of samples for each language in our subset.
Evaluation: We evaluate [^5] Evaluation prompts and other details in Appendix §A.1 our fine-tuned models along with the available Instruction fine-tuned model variants of Mistral-7B and Phi-3-small (IFT [^6] We take the Mistral-7B instruction-tuned variant from here and Phi-3-small variant from here.) on XCOPA Ponti et al. (2020), XStoryCloze Lin et al. (2022), XWinograd Muennighoff et al. (2023)Tikhonov and Ryabinin (2021), Belebele Bandarkar et al. (2023), and XQuAD Artetxe et al. (2020) using the language model evaluation harness Gao et al. (2023).
  • XCOPA: A causal commonsense reasoning dataset in 11 languages, evaluated in a 4-shot prompt setting.
  • XStoryCloze: A professionally translated version of the English StoryCloze dataset Mostafazadeh et al. (2017) in 10 languages, evaluated in a 4-shot prompt setting.
  • Belebele: A parallel reading comprehension dataset across 122 languages, with evaluation on a subset of 14 languages in a 0-shot prompt setting.
  • XQuAD: A QA dataset consisting of professional translations of a subset of SQuAD into 10 languages, evaluated in a 3-shot prompt setting due to context window limitations.
  • XWinograd: A collection of Winograd Schemas in six languages for cross-lingual commonsense reasoning, evaluated in a 0-shot setting.
4.2 Fine-Tuning Methodology
Inspired from  Longpre et al. (2023)’s strategies to instruction tune a model, we devise Language-Specific N-shot Guided Instruction fine-tuning (LANG). With this approach, we augment a training example by prepending $N$ number of samples of same language as that of the original training example randomly selected from the corresponding few shot set. This augmented training example is used for instruction tuning our models. Suppose a training example in language $l$ is a pair of $Instruction({I}_{\text{train}}^{l})$ and $Response({R}_{\text{train}}^{l})$. We need to prepend $N$ number of shots represented by $({I}_{\text{fewshot}_{1}}^{l},{R}_{\text{fewshot}_{1}}^{l})$, $({I}_{\text{fewshot}_{2}}^{l},{R}_{\text{fewshot}_{2}}^{l})$, …, $({I}_{\text{fewshot}_{N}}^{l},{R}_{\text{fewshot}_{N}}^{l})$. For every training example, $N$ is chosen with a probability $p$ as defined in Table 3. $\mathcal{T}$ is the instruction tuning templating function which takes $Instruction$ and $Response$ and transforms it to user-assistant format by adding the special tokens. Thus, the final training example is:
$(\bigoplus_{i=1}^{N}\mathcal{T}(I_{\text{fewshot}_{i}}^{l},R_{\text{fewshot}_{i}}^{l})\oplus\mathcal{T}(I_{\text{train}}^{l},R_{\text{train}}^{l}))$
where $N$ is chosen based on the probability distribution $P(N)$. $\bigoplus$ represents combining all few-shot examples, while $\oplus$ indicates concatenating the main training example with the aggregated few-shot examples. Once $N$ is selected, an equivalent number of few-shot examples is sampled uniformly at random from the few shot set. The maximum context token length for training the models is set at 8192. To ensure augmented samples fall within this range, we assign more weights to $N$=0 and $N$=1. This strategy is consistently applied when instruction tuning other datasets for both Phi-3-small and Mistral-7B base model. The average token length of samples in each datasets as per Phi-3-small model tokenizer is shown in Table 1. To assess the effectiveness of each instruction tuning dataset on an equal scale, we conducted a comparative analysis of model performance, fine-tuning each on approximately 8 billion tokens from each dataset using LANG strategy.
4.3 Hyperparameters and Training Setup
We used 5 nodes with each node containing 8 A100 GPUs with 80GB VRAM. These nodes communicated with each other using InfiniBand  [^7] https://network.nvidia.com/pdf/whitepapers/IB_Intro_WP_190.pdf. We use DeepSpeed Rasley et al. (2020) to do distributed fine-tuning over these GPUs. We use the same hyperparameters (Table 2) to fine-tune both Mistral-7B and Phi-3-small models.
5 Results
Figure 4: Performance of Mistral-7B and Phi-3-smallwhen instruction-tuned on 8B tokens across various datasets on different benchmarks.
Figure 5: Overall performance for both Mistral-7B and Phi-3-small on our dataset and fine-tuning recipe. Legend: Base_Model: Base Model, IFT: Instruction Fine-Tuned variant of the Base Model available publicly, Bactrian: Base model trained on Bactrian dataset, mAlpaca: Base Model trained on mAlpaca dataset, Aya: Base Model trained on a subset of the Aya dataset, Sphinx-0s: Base Model trained in a 0-shot fashion on the sPhinX dataset, Sphinx: Base model trained using LANG on sPhinX dataset
We evaluate the reasoning, question answering, and reading comprehension abilities of the Phi-3-small and Mistral-7B models, instruction-tuned on different multilingual datasets, using various benchmarks and find that fine-tuning on sPhinX provides an improvement of 4.2%pt [^8] A percentage point (pt) is the unit for expressing the absolute difference between two percentage values and 5%pt respectively ( language-wise results are in the Appendix §A.2). Additionally as observed in Figure 4, the sPhinX dataset significantly enhances the multilingual performance of the Phi-3-small and Mistral-7B model compared to other datasets even when finetuned on equal number of tokens. Furthermore, during instruction tuning on 8B tokens, the models encountered fewer training samples for sPhinX due to its higher average token length per sample, as illustrated in Table 1.
Due to the code-mixed nature of the instruction along with CoT reasoning explanations, a single sample of sPhinX is notably richer as compared to its counterparts from the other datasets. Consequently, even with fewer samples (keeping the number of the tokens same), models trained on sPhinX achieve better performance; thereby demonstrating the per sample efficiency of sPhinX.
6 Ablation: Improvements from LANG
To demonstrate the effectiveness of our LANG strategy, we also instruction-tuned the models on sPhinX without any pre-pended shots, referring to this as sPhinX-0s. As shown in Figure 5 (with detailed results in Appendix §A.2), models fine-tuned on sPhinX especially Mistral-7B exhibits superior performance compared to its counterparts fine-tuned on other datasets across all benchmarks. Moreover, fine-tuning both Phi-3-small and the Mistral-7B on sPhinX using the LANG strategy further boosts the performance by 3.2%pt and 10%pt respectively as compared to the base model fine-tuned without this strategy. Surprisingly, even without the LANG strategy, models instruction-tuned on sPhinX still perform better than those tuned on other datasets with LANG.
Furthermore, employing the LANG strategy leads to additional performance improvements indicating that LANG can effectively enhance the multilingual capabilities of LLMs. From the detailed results in Appendix §A.2, we observe no performance regression on high resource languages which normally occurs due to catastrophic forgetting Chang et al. (2023).
We also observe significant performance improvements in medium and low-resource languages such as Arabic, Hindi, Thai, Turkish, Tamil, and Telugu, further showcasing the effectiveness of our dataset and the LANG fine-tuning strategy (Appendix §A.2).
6.1 Regression Analysis on Standard LLM Benchmarks
It is well-studied that training on multiple languages cause regression on the performance on English language due to catastrophic forgetting Chang et al. (2023). We test this by checking for performance of the Phi-3-small model fine-tuned with sPhinX on English in the multilingual benchmarks we evaluate ((Appendix §A.2) and on popular English-only benchmarks.
We find that the Phi-3-small sPhinX model maintains its performance in English on the multilingual benchmarks and is also consistently able to maintain performance on standard English benchmarks such as MMLU (5-shot) Hendrycks et al. (2021), MedQA (2-shot) Jin et al. (2021), Arc-C (10-shot), Arc-E (10-shot) Clark et al. (2018), PiQA (5-shot) Bisk et al. (2020), WinoGrande (5-shot) Sakaguchi et al. (2021), OpenBookQA (10-shot) Mihaylov et al. (2018), BoolQ (2-shot) Clark et al. (2019) and CommonSenseQA (10-shot) Talmor et al. (2018) (Table 10). We notice some regression in the GSM-8k (8-shot, CoT) Cobbe et al. (2021) benchmark. This indicates that gains in multilingual performance caused by sPhinX do not come at the cost of regression in English performance.
7 Conclusion
In this paper, we demonstrated how instruction tuning Phi-3-small and Mistral-7B on sPhinX effectively improve their multilingual capabilities. We observed that instruction tuning the models using the sPhinX dataset leads to consistent gains on an average by 4.2%pt for Phi-3-small and 5%pt for Mistral-7B respectively. We find that the Phi-3-small model instruction tuned on sPhinX is the best performing model on multilingual benchmarks. Moreover, sPhinX exhibits greater sample efficiency and diversity compared to other multilingual instruction tuning datasets.
Additionally, we proposed a method for further enhancing the model’s performance by utilizing LANG, which supplements the training examples with $N$ samples from a few-shot set, providing the model with additional context to aid in its learning process. This further boosts the performance for both Phi-3-small and Mistral-7B by 3.2%pt and 10%pt respectively. Models instruction-tuned on sPhinX, exhibit enhanced performance even in languages they have not previously encountered during training. Finally, we evaluate the performance of the Phi-3-small model fine-tuned on sPhinX and find that it is able to maintain English performance, suggesting that gains in multilingual performance do not come at the cost of English performance while fine-tuning with sPhinX. Through our comprehensive experiments, analyses, and findings, we aim to contribute to the progress of LLMs for multilingual purposes, promoting advancements in natural language processing across a broader spectrum of languages.
8 Future Work
We have conducted all experiments using 7B base models in full parameter fine-tuning settings. It would be interesting to study the same effect by adaptive learning using LoRA Hu et al. (2022) or PEFT Mangrulkar et al. (2022). We believe that our strategies could also be effective for models with fewer parameters, leading to a notable improvement in multilingual performance.
Our LANG strategy involves utilizing $N$ examples from the same language. Other ideas that can be explored are incorporating $N$ examples from the same language script to increase the diversity of the sample. This approach could be particularly beneficial for enhancing the performance of models in low-resource languages.
Limitations
Our study has several limitations that can be considered in future research. Firstly, we conducted an extensive series of experiments, utilizing significant GPU resources and substantial time for model fine-tuning. Due to these resource-intensive processes, it may be difficult to apply our strategies to fully fine-tune a model. Besides, our study is confined to 7B models, explicitly excluding larger models. Despite this limitation, we believe our methodologies are broadly applicable for fine-tuning smaller datasets using techniques like LoRA and PEFT. Secondly, our fine-tuning dataset focuses on reasoning tasks and excludes some low-resource languages. We evaluated the models’ performance against these reasoning benchmarks. However, we did not benchmark our models on generative tasks such as summarization, nor did we evaluate models on hallucination, toxicity, or fairness.
Ethics Statement
Despite our rigorous efforts to ensure that our dataset is free from discriminatory, biased, or false information, there remains a possibility that these problems are present, particularly in multilingual contexts. Hence, it is possible that these issues might propagate to our fine-tuned models as well. We are committed to mitigate such risks and strongly advocate for the responsible use of recipes and prevent any unintended negative consequences.
Acknowledgements
We thank the Orca team at Microsoft Research for sharing the Orca dataset to implement the synthetic dataset creation recipe.
Appendix A Appendix
A.1 Prompt Templates
Figure 6 is the template for Selective Translation that was used to generate the synthetic data. Our reference dataset is in English and the {language} is the target language to generate the data in. Figure 7, 8, 9,10 and 11 are the prompts used to evaluate XQuAD, XstoryCloze, Xwinograd, XCOPA and Belebele respectively.
A.2 Detailed Results
Tables 4, 5, 6, 7 and 8 show the granular results on our models and dataset.
[…]
Please carefully convert a conversation between a human and an AI assistant from English to language. The dialogue will be presented in JSON format, where ’system’ denotes system instructions, ’human’ indicates user queries, and ’assistant’ refers to the AI’s response. You should approach this task as if the ’human’ original language is {language}. Translate the ’system’ instructions fully into {language}. For the ’human’ input, however, carefully discern which segments require translation into {language}, while leaving other parts in their original form.
For instance: 1. If the human contains a mix of languages, only translate the instruction part.
2. If the task is about language correction do not translate the target passage.
For the ’assistant’ part, generate the ’assistant’ response as you were prompted with ths newly translated system and assistant instructions. The outcome should retain the JSON format. Your response should solely contain the JSON. Do not translate the JSON keys. {"system": System text here, "human": User text here, "assistant": Assistant text here }
Figure 6: Prompt for Selective Translation using GPT-4
[…]
The task is to solve reading comprehension problems. You will be provided questions on a set of passages and you will need to provide the answer as it appears in the passage. The answer should be in the same language as the question and the passage.
Context:
{context}
Question:
{question}
Referring to the passage above, the correct answer to the given question is {answer}
Figure 7: XQuAD evaluation prompt
[…]
{input_sentence_1} {input_sentence_2}
{input_sentence_3} {input_sentence_4}
What is a possible continuation for the story given the following options?
Option1: {sentence_quiz1} Option1: {sentence_quiz2}
Figure 8: XstoryCloze evaluation prompt
[…]
Select the correct option out of option1 and option2 that will fill in the _ in the below sentence:
{sentence}
Choices:
-option1: {option1}
-option2: {option2}
Figure 9: Xwinograd evaluation prompt
[…]
The task is to perform open-domain commonsense causal reasoning. You will be provided a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise. Answer as concisely as possible in the same format as the examples below: Given this premise:
{premise}
What’s the best option?
-choice1 : {choice1}
-choice2 : {choice2}
We are looking for{if question == c̈ause%̈} a cause {else %} an effect {endif %}
Figure 10: XCOPA evaluation prompt
[…]
The task is to perform reading comprehension task. Given the following passage, query, and answer choices, output only the letter corresponding to the correct answer. Do not give me any explanations to your answer. Just a single letter corresponding to the correct answer will suffice.
Passage: {flores_passage}
Query: {question}
Choices:
A: {mc_answer1}
B: {mc_answer2}
C: {mc_answer3}
D: {mc_answer4}
Figure 11: Belebele evaluation prompt
Hyperparameter Value
Batch Size 512
Context length 8192
Learning Rate $10*e^{-5}$
Scheduler Cosine
Epochs 10
Weight Decay 0.1
Optimizer AdamW
Table 2: Hyperparameters for model fine-tuning
$N$ $p(N)$ $N$ $p(N)$
0 0.3 4 0.1
1 0.2 5 0.1
2 0.1 6 0.1
3 0.1
Table 3: Probabilities of selecting number of shots in the LANG strategy
Language en fr jp pt ru zh avg
Mistral-7B
Base Model 0.52 0.47 0.52 0.54 0.54 0.50 0.52
IFT 0.61 0.57 0.57 0.57 0.60 0.56 0.58
m-Alpaca 0.61 0.57 0.57 0.57 0.60 0.56 0.58
Aya 0.55 0.56 0.54 0.54 0.56 0.54 0.55
Bactrian 0.61 0.57 0.57 0.57 0.60 0.56 0.58
sPhinX-0s 0.75 0.65 0.68 0.67 0.66 0.65 0.68
sPhinX 0.80 0.69 0.72 0.70 0.67 0.67 0.71
Phi-3-small
Base Model 0.86 0.67 0.73 0.77 0.74 0.72 0.75
IFT 0.86 0.78 0.72 0.78 0.77 0.75 0.78
m-Alpaca 0.87 0.76 0.75 0.78 0.76 0.71 0.81
Aya 0.79 0.61 0.67 0.70 0.70 0.66 0.69
Bactrian 0.83 0.72 0.71 0.75 0.70 0.68 0.73
sPhinX-0s 0.89 0.77 0.78 missing0.79 0.81 0.76 0.80
sPhinX missing0.89 missing0.76 missing0.79 0.78 missing0.82 missing0.77 missing0.84
Table 4: Language-wise performance of instruction-tuned Mistral-7B and Phi-3-small  models evaluated on XWinograd (0-shot). Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar de el en es hi ro ru th tr vi zh avg
Mistral-7B
Base Model 0.62 0.81 0.64 0.89 0.86 0.65 0.82 0.71 0.59 0.68 0.79 0.76 0.74
IFT 0.42 0.68 0.33 0.92 0.66 0.5 0.71 0.61 0.38 0.63 0.71 0.68 0.60
m-Alpaca 0.10 0.75 0.15 0.86 0.82 0.12 0.62 0.68 0.12 0.38 0.52 0.46 0.46
Aya 0.33 0.73 0.65 0.85 0.8 0.63 0.75 0.67 0.57 0.61 0.75 0.59 0.66
Bactrian 0.67 0.76 0.26 0.85 0.86 0.74 0.77 0.71 0.59 0.69 0.77 0.65 0.69
sPhinX-0s 0.54 0.76 0.7 0.88 0.84 0.69 0.77 0.66 0.52 0.64 0.71 0.60 0.69
sPhinX 0.74 0.87 0.76 0.93 0.90 0.79 0.86 0.77 0.63 0.76 0.88 0.73 0.80
Phi-3-small
Base Model 0.68 0.90 0.77 0.93 0.91 0.61 0.84 0.80 0.55 0.73 0.86 0.69 0.78
IFT 0.71 0.88 0.73 0.92 0.91 0.64 0.84 0.80 0.44 0.70 0.67 0.76 0.75
m-Alpaca 0.55 0.92 0.74 0.96 0.94 0.68 0.87 0.85 0.50 0.73 0.88 0.66 0.77
Aya 0.61 0.89 0.84 0.94 0.93 0.80 0.89 0.82 0.73 0.83 0.92 0.79 0.83
Bactrian 0.81 0.92 0.81 0.95 0.95 0.80 0.91 0.84 0.74 0.82 0.92 0.79 0.85
sPhinX-0s 0.75 0.89 0.81 0.94 0.94 0.75 0.87 0.79 0.63 0.77 0.88 0.78 0.82
sPhinX missing0.84 missing0.93 missing0.87 missing0.96 missing0.96 missing0.81 missing0.91 missing0.86 missing0.73 missing0.84 missing0.92 missing0.81 missing0.87
Table 5: Granular results for XQuAD (3-shot) on our model. Metric: F1. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language et ht id it qu sw ta th tr vi zh en avg
Mistral-7B
Base Model 0.54 0.51 0.72 0.81 0.49 0.52 0.50 0.53 0.58 0.62 0.78 0.93 0.63
IFT 0.52 0.52 0.69 0.79 0.50 0.51 0.50 0.54 0.57 0.63 0.75 0.90 0.62
m-Alpaca 0.51 0.50 0.52 0.63 0.50 0.50 0.50 0.51 0.51 0.49 0.65 0.74 0.55
Aya 0.57 0.54 0.64 0.67 0.53 0.56 0.57 0.62 0.56 0.61 0.64 0.78 0.61
Bactrian 0.52 0.50 0.53 0.60 0.49 0.51 0.50 0.51 0.51 0.52 0.52 0.71 0.54
sPhinX-0s 0.54 0.5 0.58 0.63 0.51 0.55 0.52 0.52 0.54 0.57 0.64 0.8 0.58
sPhinX missing0.64 0.54 0.73 0.80 0.53 0.61 0.59 0.63 0.67 0.66 0.80 0.91 0.68
Phi-3-small
Base Model 0.55 0.51 0.80 0.93 0.52 0.54 0.46 0.56 0.61 0.66 0.86 0.98 0.64
IFT 0.55 0.57 0.81 0.93 0.53 0.58 0.48 0.60 0.62 0.69 0.88 0.96 0.68
m-Alpaca 0.53 0.54 0.80 0.92 0.49 0.54 0.51 0.59 0.64 0.68 0.87 0.99 0.68
Aya 0.60 0.55 0.72 0.83 0.52 0.55 0.52 0.62 0.59 0.69 0.75 0.89 0.65
Bactrian 0.62 0.56 0.83 0.91 0.54 0.62 0.52 0.66 0.65 0.71 0.86 0.99 0.71
sPhinX-0s 0.59 0.58 0.85 0.94 0.50 0.60 0.54 0.63 0.69 missing0.72 0.89 0.96 0.71
sPhinX 0.59 missing0.60 missing0.85 missing0.94 missing0.52 missing0.57 missing0.58 missing0.68 missing0.68 0.71 missing0.90 missing0.98 missing0.72
Table 6: Granular results for XCOPA (4-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar en es eu hi id my ru sw te zh avg
Mistral-7B
Base Model 0.65 0.89 0.83 0.56 0.62 0.76 0.52 0.81 0.56 0.52 0.80 0.68
IFT 0.70 0.96 0.92 0.54 0.69 0.79 0.57 0.90 0.58 0.54 0.88 0.73
m-Alpaca 0.53 0.73 0.7 0.51 0.51 0.57 0.50 0.66 0.52 0.52 0.71 0.59
Aya 0.64 0.86 0.81 0.56 0.71 0.73 0.60 0.82 0.67 0.60 0.81 0.71
Bactrian 0.69 0.82 0.74 0.52 0.59 0.76 0.54 0.73 0.62 0.61 0.76 0.67
sPhinX-0s 0.57 0.66 0.64 0.47 0.56 0.61 0.50 0.62 0.56 0.52 0.69 0.58
sPhinX 0.83 0.96 0.94 0.57 missing0.84 0.87 missing0.61 0.91 missing0.80 missing0.69 0.94 0.81
Phi-3-small
Base Model 0.80 0.98 0.96 0.61 0.72 0.92 0.53 0.96 0.61 0.55 0.94 0.78
IFT 0.81 0.98 0.96 0.61 0.75 0.92 0.56 0.96 0.61 0.53 0.94 0.79
m-Alpaca 0.81 0.99 0.98 0.58 0.76 0.93 0.52 0.97 0.64 0.54 0.96 0.79
Aya 0.77 0.98 0.97 0.57 0.77 0.93 0.53 0.96 0.74 0.56 0.94 0.79
Bactrian 0.83 0.98 0.98 0.61 0.83 0.94 0.54 0.97 0.79 0.63 0.94 0.82
sPhinX-0s 0.84 0.98 0.97 missing0.64 0.77 0.95 0.52 0.96 0.74 0.57 0.95 0.81
sPhinX missing0.86 missing0.99 missing0.99 0.61 0.82 missing0.96 0.54 missing0.98 0.74 0.61 missing0.97 missing0.82
Table 7: Granular results for XStoryCloze (4-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar de es en fi fr hi it jp ko ta te vi zh avg
Mistral-7B
Base Model 0.25 0.23 0.23 0.24 0.23 0.23 0.26 0.24 0.26 0.23 0.23 0.25 0.26 0.25 0.24
IFT 0.32 0.60 0.62 0.74 0.36 0.62 0.32 0.61 0.43 0.47 0.27 0.27 0.39 0.58 0.47
m-Alpaca 0.32 0.50 0.53 0.56 0.45 0.51 0.27 0.51 0.40 0.41 0.26 0.26 0.33 0.48 0.41
Aya 0.34 0.43 0.43 0.48 0.38 0.47 0.35 0.44 0.4 0.36 0.27 0.25 0.37 0.42 0.38
Bactrian 0.24 0.27 0.25 0.25 0.26 0.27 0.24 0.28 0.26 0.26 0.23 0.23 0.34 0.28 0.26
sPhinX-0s 0.64 0.75 0.75 0.82 0.66 0.79 0.53 0.73 0.69 0.66 0.48 0.44 0.66 0.75 0.67
sPhinX 0.69 0.80 0.69 0.87 0.71 0.82 missing0.60 0.79 0.73 0.73 missing0.56 missing0.48 0.70 0.80 0.71
Phi-3-small
Base Model 0.54 0.87 0.85 0.92 0.58 0.86 0.41 0.86 0.70 0.58 0.26 0.30 0.62 0.82 0.65
IFT 0.63 0.89 0.88 0.93 0.63 0.88 0.48 0.88 0.77 0.68 0.32 0.32 0.68 0.85 0.70
m-Alpaca 0.65 0.92 0.90 0.94 0.74 0.91 0.54 0.90 0.80 0.70 0.47 0.45 0.72 0.84 0.75
Aya 0.58 0.86 0.85 0.91 0.65 0.87 0.50 0.86 0.76 0.67 0.37 0.35 0.69 0.84 0.70
Bactrian 0.67 0.88 0.88 0.92 0.70 0.88 0.51 0.86 0.77 0.70 0.37 0.37 0.74 0.86 0.72
sPhinX-0s 0.73 0.91 0.90 0.93 0.75 0.92 0.57 0.91 0.82 missing0.82 0.45 0.40 0.76 0.89 0.77
sPhinX missing0.74 missing0.93 missing0.91 missing0.94 missing0.77 missing0.93 0.58 missing0.92 missing0.84 0.76 missing0.46 0.40 missing0.78 missing0.89 missing0.79
Table 8: Granular results for Belebele (0-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Model XC XS XW XQ BL
Mistral-7B
Base Model 0.63 0.68 0.52 0.74 0.24
IFT 0.62 0.73 0.54 0.60 0.47
m-Alpaca 0.55 0.59 0.51 0.46 0.41
Aya 0.68 0.71 0.54 0.66 0.38
Bactrian 0.54 0.67 0.54 0.69 0.26
sPhinX-0s 0.58 0.58 0.68 0.69 0.67
sPhinX 0.68 0.81 0.71 0.80 0.71
Phi-3-small
Base Model 0.64 0.78 0.75 0.78 0.65
IFT 0.68 0.79 0.78 0.75 0.70
m-Alpaca 0.68 0.79 0.81 0.77 0.75
Aya 0.65 0.79 0.69 0.83 0.72
Bactrian 0.71 0.82 0.73 0.85 0.77
sPhinX-0s 0.71 0.81 0.80 0.82 0.79
sPhinX 0.72 0.84 0.87 0.87 0.79
Table 9: Performance of Mistral-7B and Phi-3-small instruction-tuned for 10000 training steps on various datasets. Abbreviations: XC - XCOPA, XS - XStoryCloze, XW - XWinograd, XQ - XQuAD, BL - Belebele. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Benchmarks Base Model sPhinX
MMLU
(5-shot)
0.76 0.75
HellaSwag
(5-shot)
0.81 0.83
GSM-8k
(8-shot, CoT)
0.85 0.77
MedQA
(2-shot)
0.64 0.66
Arc-C
(10-shot)
0.90 0.90
Arc-E
(10-shot)
0.97 0.97
PIQA
(5-shot)
0.84 0.89
WinoGrande
(5-shot)
0.77 0.82
OpenBookQA
(10-shot)
0.86 0.88
BoolQ
(2-shot)
0.82 0.87
CommonSenseQA
(10-shot)
0.80 0.81
Table 10: Performance of the Phi-3-small base model and the sPhinX tuned model on standard English LLM benchmarks.
High-Resource (100k)
Spanish, Chinese Simplified, Japanese
French, German, Portuguese, Italian
Mid-Resource (50k)
Dutch, Swedish, Danish
Finnish, Russian, Norwegian
Korean, Chinese Traditional, Polish
Turkish, Arabic, Hebrew
Portuguese, Czech, Hungarian
Low-Resource (25k)
Indonesian, Thai, Greek
Slovak, Vietnamese, Slovenian
Croatian, Romanian, Lithuanian
Bulgarian, Serbian, Latvian
Ukranian, Estonian, Hindi
Burmese, Bengali, Afrikaan
Punjabi, Welsh, Icelandic
Marathi, Swahili, Nepali
Urdu, Telugu, Malayalam
Russian, Tamil, Oriya
Table 11: Language distribution and samples across three tiers
Code Languages Script Data
af Afrikaan Latin 20206
ar Arabic Arabic 26803
bn Bengali Bengal 20165
bg Bulgarian Cyrillic 17300
my Burmese Burmese 12123
zh-Hans Chinese_Simplified Han 100650
zh-Hant Chinese_Traditional Hant 32363
hr Croatian Latin 17340
cs Czech Latin 32711
da Danish Latin 36348
nl Dutch Latin 36586
en English Latin 199900
et Estonian Latin 17207
fi Finnish Latin 33622
fr French Latin 100337
de German Latin 100265
el Greek Greek 17317
he Hebrew Hebrew 24483
hi Hindi Devanagari 20240
hu Hungarian Latin 31999
is Icelandic Latin 20164
id Indonesian Latin 17297
it Italian Latin 85175
jp Japanese Japanese 98366
ko Korean Hangul 30890
lv Latvian Latin 17247
lt Lithuanian Latin 17232
ml Malayalam Malayalam 19817
mr Marathi Devanagari 20069
ne Nepali Devanagari 20092
nb Norwegian Latin 36811
or Oriya Oriya 19153
pl Polish Latin 34711
pt Portuguese Latin 37229
pa Punjabi Gurmukhi 20026
ro Romanian Latin 17149
ru Russian Cyrillic 20108
sr Serbian Latin 17165
sk Slovak Latin 17255
sl Slovenian Latin 17300
es Spanish Latin 100351
sw Swahili Latin 20170
sv Swedish Latin 36533
ta Tamil Tamil 19807
te Telugu Telugu 19947
th Thai Thai 17322
tr Turkish Latin 34405
uk Ukrainian Cyrillic 17282
ur Urdu Perso-Arabic 20162
vi Vietnamese Latin 17358
cy Welsh Latin 20207
Table 12: Language Distribution in Sphinx Dataset

### v4

@@
 Abstract
-Despite the remarkable success of LLMs in English, there is a significant gap in performance in non-English languages.
-In order to address this, we introduce a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX, which is created by selectively translating instruction response pairs from English into 50 languages.
-We test the effectiveness of sPhinX by using it to fine-tune two state-of-the-art models, Phi-3-small and Mistral-7B and then evaluating them across a comprehensive suite of multilingual benchmarks that test reasoning, question answering, and reading comprehension.
-Our results show that Phi-3-small and Mistral-7B fine-tuned with sPhinX perform better on an average by 4.2%pt and 5%pt respectively as compared to the baselines.
-We also devise a strategy to incorporate N-shot examples in each fine-tuning sample which further boosts the performance of these models by 3%pt and 10%pt respectively.
-sPhinX outperforms other multilingual instruction tuning datasets on the same benchmarks along with being sample efficient and diverse, thereby reducing dataset creation costs.
-Additionally, instruction tuning with sPhinX does not lead to regression on most standard LLM benchmarks.
+Despite the remarkable success of large language models (LLMs) in English, a significant performance gap remains in non-English languages.
+To address this, we introduce a novel approach for strategically constructing a multilingual synthetic instruction tuning dataset, sPhinX .
+Unlike prior methods that directly translate fixed instruction-response pairs, sPhinX enhances diversity by selectively augmenting English instruction-response pairs with multilingual translations.
+Additionally, we propose LANGIT, a novel N-shot guided fine-tuning strategy, which further enhances model performance by incorporating contextually relevant examples in each training sample.
+Our ablation study shows that our approach enhances the multilingual capabilities of Mistral-7B and Phi-3-small  improving performance by an average of 39.8% and 11.2%, respectively, across multilingual benchmarks in reasoning, question answering, reading comprehension, and machine translation.
+Moreover, sPhinX maintains strong performance on English LLM benchmarks while exhibiting minimal to no catastrophic forgetting, even when trained on 51 languages.
 1 Introduction
-Large Language Models (LLMs) have been shown to perform several tasks exceptionally well in English, however, performance in some non-English languages still lags behind Ahuja et al.
+Large Language Models (LLMs) have demonstrated exceptional performance across various tasks in English.
+However, their performance in some non-English languages remains comparatively limited Ahuja et al.
 (2023); Asai et al.
 (2024).
-Further, the gap between the performance of Large Language Models (LLMs) and Small Language Models (SLMs) is more pronounced Ahuja et al.
-(2024) on
-non-English languages.
+Further, the gap between the performance of Large Language Models (LLMs) and Small Language Models (SLMs) is more pronounced (Ahuja et al., 2024) in non-English languages.
 Cui et al.
-(2023) and Balachandran (2023) follow the approach of training the models further on datasets for specific languages, however this can lead to catastrophic forgetting and hurt the performance on English Zhao et al.
+(2023) and Balachandran (2023) utilize the method of fine-tuning models on datasets focused on particular languages.
+However, this can lead to catastrophic forgetting, which may negatively impact performance in English Zhao et al.
+(2024); Aggarwal et al.
 (2024).
-Several techniques have been proposed to bridge this gap, such as incorporating better pre-training data in various languages and enhancing base tokenizers Xu et al.
+Few techniques have been proposed to bridge this gap, such as incorporating better pre-training data in various languages and improving base tokenizers Xu et al.
 (2024); Dagan et al.
 (2024).
-However, most of these changes need to be implemented at the pre-training stage which demands extensive data and computational resources, making it practically unfeasible in many scenarios Brown et al.
+However, most of these changes need to be implemented in the pre-training stage, which demands extensive data and computational resources, making it practically unfeasible in many scenarios Brown et al.
 (2020).
 Consequently, the most well-studied technique involves fine-tuning models for specific languages and tasks.
-Instruction fine-tuning (IFT) has become a popular technique for enhancing the performance of language models in specific languages.
-This method combines the benefits of both the pretrain–fine-tune and prompting paradigms Wei et al.
+Instruction fine-tuning (IFT) has become a popular technique to enhance the performance of language models in specific languages.
+This method combines the benefits of both the pre-training, fine-tuning, and prompting paradigms Wei et al.
 (2021).
-Sample diversity is crucial for instruction tuning in multilingual datasets.
-Many recent datasets are created by translating English content into other languages or using self-instruct techniques from seed prompts Li et al.
+Sample diversity is essential for effective instruction tuning in multilingual datasets.
+Many recent datasets have been generated by translating English content into other languages or by employing self-instruct techniques based on seed prompts Li et al.
 (2023); Taori et al.
-(2023), both of which can limit diversity.
-Machine translation can lead to loss of semantic information Baroni and Bernardini (2006), while self-instruct methods may produce repetitive and homogeneous samples Wang et al.
 (2023).
-This underscores the importance of having a dataset with diverse set of samples.
-Aggarwal et al.
-(2024) evaluate several models fine-tuned using Parameter Efficient fine-tuning and find that there is a gain in multilingual performance across several SLMs for many low-resource languages, with some high-resource languages performing worse after fine-tuning.
-However, the performance gains often do not match the performance of larger models, such as GPT-4 and Gemini and can be inconsistent across languages.
-Hence, there is a need to study instruction tuning for better multilingual performance in SLMs.
+However, both methods can limit diversity.
+Machine translation may result in the loss of semantic nuance (Baroni and Bernardini, 2006), while self-instruct approaches often yield repetitive and homogeneous samples Wang et al.
+(2022).
+This highlights the critical need for datasets that encompass a wide range of diverse samples.
 In this paper, we present a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX.
-It comprises 1.8M instruction-response pairs in 51 languages, derived by selectively translating the Orca instruction tuning dataset Mukherjee et al.
-(2023) using GPT-4 Achiam et al.
+It comprises 1.8M instruction-response pairs in 51 languages, derived by augmenting the Orca instruction tuning dataset samplesMukherjee et al.
+(2023) through Selective Translated Augmentation using GPT-4 Achiam et al.
 (2023).
-We assess the effectiveness of sPhinX by fine-tuning two models — Phi-3-small and Mistral-7B — across a range of evaluation benchmarks that test various language model capabilities like reasoning, question answering, and reading comprehension.
+We assess the effectiveness of sPhinX by fine-tuning two models — Phi-3-small and Mistral-7B — across a range of evaluation benchmarks that test various language model capabilities across discriminative and generative tasks.
 We compare models fine-tuned on sPhinX with those using other synthetic multilingual instruction tuning datasets like Aya Üstün et al.
@@
 (2023), and Bactrian Li et al.
-(2023) and observe significant performance gains.
+(2023) and observe significant performance gains across languages.
+We also compare our proposed translation strategy with translating the entire instruction using Azure Translator API, as is done with the popular multilingual synthetic IFT datasets to demonstrate the efficacy of our approach.
 The contributions of this paper are as follows:
-  • We introduce a novel approach to generate synthetic data for multilingual instruction tuning by selectively translating the Orca dataset with the assistance of GPT-4 (§3.1)
-  • We devise LAnguage-Specific N-shot Guided Instruction fine-tuning (LANG) strategy for enhancing the multilingual capabilities of LLMs (§4.2)
-  • We conduct an extensive number of experiments on different multilingual instruction tuning datasets to test generalizability in multilingual settings (§5)
-  • We also conduct an in-depth analysis where we find sPhinX to be more sample efficient and diverse across languages (§3.3)
+  • We introduce a novel approach to generate synthetic data for multilingual instruction tuning by Selective Translated Augmentation of the Orca dataset with the assistance of GPT-4 (§3.1)
+  • We devise LAnguage-Specific N-shot Guided Instruction Tuning (LANGIT) strategy for enhancing the multilingual capabilities of LLMs (§4)
+  • We also conduct extensive instruction tuning experiments on various multilingual instruction tuning datasets to evaluate generalizability in multilingual settings (§6).
+  • We plan to release a subset of the augmented dataset by applying our strategy to the OpenOrca  [^1] https://huggingface.co/datasets/Open-Orca/OpenOrca dataset (Lian et al., 2023) (Open-sPhinX) as well.
 2 Related Work
-2.1 Multilingual LLMs
-Recently, there has been there has been interest in developing and improving SLMs such as LLaMA Touvron et al.
-(2023), Mistral Jiang et al.
-(2023), Phi3 Abdin et al.
-(2024), and Gemma Team et al.
-(2024).
-In a relatively brief period, researchers have developed numerous monolingual and multilingual versions of these foundational models, either by pre-training from scratch on specific languages or groups of languages, or by further fine-tuning on them Nguyen et al.
-(2023); Gala et al.
-(2024); Balachandran (2023); Cui et al.
-(2023); Zhang et al.
-(2023); Qin et al.
-(2024).
-Prior work has shown that smaller models show promising results in high-resource languages but perform worse on low-resource ones Ahuja et al.
-(2024) compared to larger models.
-To address this, researchers have proposed using multilingual corpora for pre-training, expanding the vocabulary Wang et al.
-(2019) and continual training Liu et al.
-(2023), though this increases training costs significantly.
-fine-tuning for specific tasks, such as translation, has been explored but often struggles with generalization  Mishra et al.
-(2021).
-In contrast, instruction tuning has shown to improve generalization to unseen tasks, enhancing the models’ ability to understand and respond to human instructions  Mishra et al.
-(2021); Ouyang et al.
-(2022).
-This makes the models more versatile and capable of handling a wide variety of tasks.
 2.1 Multilingual Instruction fine-tuning
@@
 (2022).
-Currently, the approach has shifted on fine-tuning these models using a wide-variety of tasks Longpre et al.
+Currently, the approach has shifted to fine-tuning these models using a wide variety of tasks Longpre et al.
 (2023); Ouyang et al.
@@
 (2022) and mT0 Muennighoff et al.
-(2022) makes significant strides in improving the multilingual performance of decoder-based models Ahuja et al.
+(2022) make significant strides in improving the multilingual performance of decoder-based models Ahuja et al.
 (2023).
-There have been multiple multilingual instruction datasets and models proposed such as Bactrian Li et al.
-(2023), AYA Üstün et al.
-(2024), polyLM Wei et al.
-(2023b) after BLOOMZ and mT0.
-However, these models still do not perform as well as English on other languages, with the gap being particularly large for low-resource languages and languages written in scripts other than the Latin script  Ruder et al.
+There have been multiple multilingual instruction datasets and models proposed such as Bactrian (Li et al., 2023), AYA (Üstün et al., 2024), polyLM (Wei et al., 2023b) after BLOOMZ and mT0 Muennighoff et al.
+(2023).
+However, these models still do not perform as well as English in other languages, with the gap being huge for low-resource languages and languages written in scripts other than the Latin script  Ruder et al.
 (2021); Ahuja et al.
@@
 In this work, we aim to narrow the performance gap by introducing a strategy for creating datasets for multilingual instruction tuning and recipes for fine-tuning, which we will discuss in the following sections.
-Figure 1: The above figure describes translating using a Translation API vs Selective Translation
+Figure 1: The figure above illustrates pipelines for sPhinX data creation using Selective Translated Augmentation and Multilingual Instruction Tuning using LANGIT strategy.
+Figure 2: The figure compares the Translation API with Selective Translated Augmentation.
+The Translation API translates the entire input into Russian, while the Selective strategy localizes only necessary components.
+Here, system and user prompts are translated, but the input question and assistant’s response remain in the original language, preserving structure and intent.
 2.2 Multilingual Synthetic Data Generation
 Most instruction-tuning datasets across multiple languages typically focus on general tasks rather than specific reasoning capabilities.
-While datasets like Orca Mukherjee et al.
+Although datasets like Orca Mukherjee et al.
 (2023) and Orca 2 Mitra et al.
@@
 (2023) translates Alpaca and Dolly instructions into 52 languages using the Google Translator API and generates outputs with GPT-3.5 turbo.
-However, these translated datasets often struggle to encode semantic information effectively Baroni and Bernardini (2006).
+Another example is of this work  Lai et al.
+(2024), which also utilizes Google Translation API for translating their source datasets.
 Our dataset approach aims to tackle these challenges by selectively translating only the essential portions of multilingual inputs.
 This strategy not only preserves semantic information but also accommodates diverse linguistic contexts, thereby enhancing the overall quality and applicability of instruction-tuning datasets across languages.
+In our work, we also show the pitfalls of training with data generated only using the Translation APIs such Google Translate or Azure Translate.
 3 sPhinX Dataset
-In this section, we describe our dataset construction methodology (§3.1), dataset filtering and cleaning pipelines (§3.2), and exploratory analysis (§3.3) for determining diversity of our data in comparison to other instruction tuning datasets.
-Figure 2: Some examples of input queries and its counterpart existing in the hindi version of the Multialpaca dataset and if it was generated using the Selective Translation strategy
+In this section, we describe our dataset construction methodology (§3.1), dataset filtering, and cleaning pipelines (§3.2).
 3.1 Dataset Construction
-Mukherjee et al.
-(2023) illustrate how smaller models can replicate the reasoning processes of LLMs and learn from detailed signals like as explanation traces, step-by-step thought processes, and other intricate instructions obtained from data annotated by ChatGPT and GPT-4111GPT-4 inference hyper-parameters in Azure OpenAI interface set as: temperature=0.0, top_p=1, frequency_penalty=0, presence_penalty=0, stop=[”¡—im_end—¿”, ”¡—im_start|¿”].
-Their dataset consists of 5M ChatGPT and 1M GPT-4 generated instruction-response pairs.
-Inspired by this work, we utilized 1M GPT-4 generated instruction-response pairs of their dataset and construct our dataset along similar lines by selectively translating them into 50 different languages with the help of GPT-4.
+Inspired by (Mukherjee et al., 2023)’s work, we utilized the 1M GPT-4 generated instruction-response pairs from Orca and constructed our own dataset along similar lines using Selective Translated Augmentation into 50 different languages with the help of GPT-4222GPT-4 inference hyper-parameters in Azure OpenAI interface set as: temperature=0.0.
 We categorize them into three groups: high-resource, mid-resource, and low-resource languages as outlined in Table 7.
-For high resource languages, we randomly sample 100k instruction-response pairs from the Orca 1M dataset and generate the responses from GPT-4 using Selective Translation as shown in Figure 1.
-Similarly, we leverage the same strategy for medium and low resource languages by sampling 50k and 25k pairs respectively.
-Although GPT-4 performs competitively with commercial translation systems (Google Translate & Bing Translate) it still lags behind on medium and low resource languages Jiao et al.
+For high-resource languages, we randomly sample 100k instruction-response pairs from the Orca 1M dataset and generate the responses from GPT-4 with Selective Translated Augmentation as shown in Figure 2.
+Similarly, we leverage the same strategy for medium and low-resource languages by sampling 50k and 25k pairs respectively.
+Although GPT-4 performs competitively with commercial translation systems (Google Translate & Bing Translate) it still lags on medium and low resource languages Jiao et al.
 (2023); Hendy et al.
 (2023).
-Furthermore, as highlighted in Chang et al.
-(2023), fine-tuning with a large set of samples from medium and low-resource languages might lead to catastrophic forgetting of high-resource languages.
+Furthermore, as highlighted in (Chang et al., 2023; Lin et al., 2023; Xia et al., 2024), fine-tuning with a large set of samples from medium and low-resource languages might lead to catastrophic forgetting of high-resource languages.
 Therefore, we deliberately create fewer samples for medium and low-resource languages than for high-resource ones.
-A fundamental problem with using an off-the-shelf translation API is the lack of semantic and task awareness, in addition to translationese Baroni and Bernardini (2006)), which can result in poor quality training data.
+Besides, Shaham et al.
+(2024) also demonstrates that a small number of multilingual training samples is sufficient to significantly boost multilingual performance, validating our approach of using fewer samples from medium- and low-resource languages.
+A fundamental problem with using an off-the-shelf translation API is the lack of semantic and task awareness, in addition to translationese Baroni and Bernardini (2006), which can result in poor quality training data.
 Consider for example the task of Machine Translation as part of the instruction, wherein the language of the source sentence should be retained.
-However, an off-the-shelf API, without task awareness, would translate it, resulting in an ill defined instruction.
-To mitigate this issue, we prompt GPT-4 to selectively translate the instructions, so that the tasks are translated into the appropriate language without changing the semantic meaning.
+However, an off-the-shelf API, without task awareness, would translate it, resulting in an ambiguous instruction.
+To mitigate this issue, we used GPT-4 to augment the instructions using Selective Translated Augmentation, so that task-specific components of instruction responses are translated into the appropriate language without changing the semantic meaning.
 Figure 12 illustrates this with concrete examples.
 The first example demonstrates the aforementioned translation inconsistency issue for an instruction asking for a French equivalent of an English phrase.
-The second example demonstrates the consequence of direct translations in the MultiAlpaca dataset: wherein the translation of the task input results in the task being ill-defined based on the instructions.
-As demonstrated, our proposed Selective Translation method is able to keep the semantic information of the task intact while translating the instructions.
+The second example demonstrates the consequence of direct translations in the m-Alpaca dataset: wherein the translation of the task input results in the task being ill-defined based on the instructions.
+As demonstrated, our proposed Selective Translated Augmentation method is able to keep the semantic information of the task intact while translating the instructions.
 For the exact prompt, please refer to Figure 4 in the Appendix.
-3.2 Dataset Filtering and Cleaning
-Input: listOfSentences: list of strings
-1 Function dataFilter(listOfSentences):
-2       englishWords $\leftarrow$ set of English words from NLTK;
-3       foreach sentence in listOfSentences do
-4             cleanedSentence $\leftarrow$ replace all punctuations, digits, and single characters with a single space;
-5             cleanedSentence $\leftarrow$ replace all sequences of whitespace with a single space;
-6             wordCount $\leftarrow$ 0;
-7             foreach word in cleanedSentence do
-8                   if word.lower() in englishWords then
-9                         wordCount $\leftarrow$ wordCount + 1;
-10
-11
-12
-13            content $\leftarrow$ wordCount / len(cleanedSentence);
-14             if content $>$ 0.90 then
-15                   remove sentence from listOfSentences
-16
-17
-Algorithm 1 Data Filtering Algorithm
-Figure 3: t-SNE Visualisation of 1000 samples equally distributed in 10 languages
-After generating these translations, we analyze how many instruction-assistant pairs were unsatisfactorily translated, such as incomplete or missing translations, by identifying the English content within the instances.
-We use the NLTK [^2] https://www.nltk.org/ library to identify the fraction of English words in the sample.
-The NLTK corpus contains approximately 240,000 English words, making it suitable for our use case.
-We manually analyzed various proportions of English content in samples of each language and found that the samples were satisfactory when the English content was within 90% for all languages.
-If the English content exceeded 90% in a language that uses a different script, those samples were removed from our dataset as shown in our data filtering algorithm here.
-Finally, we manually inspected a small sample of the dataset to ensure the quality of the translations, and found them to be of good quality.
-After filtering, the final dataset comprised of 1.8 million samples in 51 languages, divided into three subsets: Train, Test, and Few-shot.
-We partitioned each language’s dataset to ensure the Test and Few-shot sets contained 2000 and 1000 samples respectively, with the Train set consisting the remaining data.
-This approach guarantees consistent distributions across languages in the Test and Few-shot sets, ensuring equitable representation regardless of training distribution.
-The train, test, and few-shot sets are in the ratio of 92 : 5.3 : 2.7.
-3.3 Data Analysis
-A primary issue that we hypothesize (and find preliminary evidence for from basic analysis of prior Multilingual IFT datasets) is the lack of diversity of samples.
-In order to validate this issue, we use t-Distributed Stochastic Neighbor Embedding (t-SNE) Cai and Ma (2022) to visualize the structure of both prior multilingual IFT datasets as well as sPhinX.
-We use the model paraphrase-multilingual-MiniLM-L12-v2 from huggingface’s sentence-transformers Reimers and Gurevych (2020), Reimers and Gurevych (2019) library [^3] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 to generate the embeddings.
-The hyperparameters were set as follows: learning rate at 200, perplexity at 30, and number of iterations (n_iter) at 1000.
-Concretely, we use 1000 samples equally distributed in 10 languages.
-Figure 3 demonstrates that sPhinX demonstrates greater diversity compared to prior datasets.
-Particularly, Bactrian and Multilingual Alpaca appear less diverse, which we hypothesize might be due to their generation method involving translations of the same set of task limited English samples into different languages.
-By virtue of using a more diverse seed set (Mukherjee et al., 2023), we circumvent this issue by design.
-Furthermore, one notable differentiating factor of sPhinX is the substantial amount of code-switched data the inherent nature of Selective Translation.
-This in turn enhances the diversity of samples across multiple languages.
-Finally, we also observe the Aya dataset to be more diverse than other prior counterparts, though it still shows less diversity across task instructions.
-Dataset Average Token
-Length/Sample
-Aya 2240
-Bactrian 2465
-Multilingual Alpaca 1620
-sPhinX-0s 544
-sPhinX 3100
-Table 1: Average Token Length in each dataset
+3.2 Dataset Filtering and Quality Assessment
+After creating the dataset, we filtered out samples where GPT-4 failed to generate a response.
+The final dataset comprised 1.8 million samples in 51 languages(Table: 15), divided into three subsets: Train, Validation, and Few-shot.
+Each language’s dataset was partitioned to ensure that the Validation and Few-shot sets contained 2,000 and 1,000 samples, respectively, while the Train set included the remaining data.
+This approach guarantees consistent distributions across languages in the Validation and Few-shot sets, ensuring equitable representation regardless of the training distribution.
+The final split ratio for Train, Validation, and Few-shot sets was 92:5.3:2.7.
+We also conducted a small-scale quality assessment of the generated data for languages such as Bengali, Hindi, German, Turkish, and Tamil.
+The researchers and engineers in our organization, who are native speakers of these languages, evaluated the data on the basis of coherence, fluency, and information retention.
+Our findings indicate that the generated dataset is moderate to high quality.
+3.3 Sample Diversity in sPhinX
+Unlike prior multilingual datasets such as Bactrian and m-Alpaca, which translate a fixed set of instruction-response pairs into multiple languages, sPhinX ensures diversity by sampling unique subsets of instruction-response pairs for each language.
+For instance, Bactrian is constructed from 67k English instruction-response pairs (Alpaca + Dolly) and translated into 52 languages, resulting in identical samples across all languages.
+In contrast, sPhinX samples from 1M GPT-4-generated instruction-response pairs, ensuring that no two languages share the exact same subset.
+Mathematically, the probability that all samples in one language dataset $A$ are also in another language dataset $B$, when sampled without replacement from a larger dataset $D$, is given by:
+$P(\text{all A in B})\approx\left(\frac{m}{N}\right)^{n}=\left(\frac{100{,}000}{1{,}000{,}000}\right)^{20{,}000}$
+where:
+  • $N=1{,}000{,}000$ (Total samples in sPhinX),
+  • $n=20{,}000$ (Samples in language $A$),
+  • $m=100{,}000$ (Samples in language $B$).
+Since the exponential term results in an extremely small probability, this confirms that no two languages have identical instruction-response sets in sPhinX.
+To further enhance diversity, we apply Selective Translated Augmentation, translating 10% of samples for high-resource languages, 5% for mid-resource languages, and 2.5% for low-resource languages.
+This ensures that translated content varies across languages, preventing uniformity.
+Additionally, code-switching naturally emerges from this augmentation process, further increasing linguistic diversity.
+Compared to Aya, which exhibits moderate variation across task instructions, sPhinX introduces greater sample diversity by leveraging a larger and more heterogeneous seed set (Mukherjee et al., 2023) and selective augmentation strategy.
+Exploring code-switching phenomena would be an interesting task in these synthetically generated datasets, but currently that is out-of-scope of for this work.
+4 LANGIT
+Following Instruction Tuning strategies of  (Longpre et al., 2023) and also taking inspiration from (Min et al., 2022), we devise Language-Specific N-shot Guided Instruction fine-tuning (LANGIT) Figure: 1.
+This method aims to improve the model’s ability to follow instructions by augmenting training examples with additional context from a set of few-shot examples in the same language.
+This added context helps guide the model, enabling it to learn more effectively from the provided examples.
+For each training example, we begin by sampling a number of few-shot examples, which are instruction-response pairs in the same language.
+The number of few-shot examples $N$ is determined probabilistically, with a 30% chance of selecting no few-shot examples, a 20% chance of selecting one, and gradually lower probabilities for higher numbers of few-shots.
+The maximum number of few-shot examples we sample is six, due to constraints imposed by the model’s context length (8192 tokens) and the typically higher tokenization length in languages other than English.
+Once the number of few-shot examples is determined, they are prepended to the main training example, forming an augmented input.
+This augmented input is then fed into the model for instruction tuning.
+The purpose of this approach is to expose the model to additional examples of different tasks, helping it generalize better to new tasks in the same language.
+We performed experiments to analyze how the model performs on each dataset when fine-tuned using the LANGIT strategy detailed in the next section (§6).
+Additionally, we fine-tuned the models on the sPhinX dataset without using LANGIT to provide a baseline for comparison.
+To assess the effectiveness of each instruction-tuning dataset on an equal scale, we conducted a comparative analysis of model performance on different benchmarks, fine-tuning each model on approximately 8 billion tokens per dataset using the LANG strategy.
+This fine-tuning strategy is consistently applied across datasets for both the Phi-3-small and Mistral-7B base models.
+A comparison of token lengths across different datasets is provided in Table 6, showing the average token lengths as tokenized by the Phi-3-small model.
 5 Experiments
 5.1 Setup
-Base Models: We use Mistral-7B  [^4] We specifically use the v1.0 base model from here.
-Jiang et al.
-(2023) and Phi-3-small Abdin et al.
-(2024) base model variants and fine-tune them on different datasets based to improve multilingual performance.
-Datasets: Apart from the sPhinX dataset, we use Bactrian Li et al. (2023), MultiAlpaca Wei et al. (2023b) and Aya Singh et al. (2024) instruction datasets for comparative evaluation §3.1.
-  • Bactrian Li et al.
-(2023) is a machine translated dataset of the original alpaca-52k Taori et al.
-(2023) and dolly-15k Conover et al.
-(2023) datasets into 52 languages.
-The instructions for this dataset were translated using a Translation API and then GPT-3.5-Turbo was prompted to generate outputs.
-We fine-tune our models on the complete dataset consisting of 3.4M instances.
-  • MultiAlpaca Wei et al.
-(2023b) is a self instruct dataset which follows the same approach as the English-only Alpaca by translating the seed instructions to 11 languages and then using GPT-3.5.-Turbo to generate responses.
-We fine-tune our models on the complete set of the dataset consisting of 500k datapoints.
-  • Aya Singh et al.
-(2024) consists of human-curated prompt-completion pairs in 65 languages called the Aya Dataset.
-It also consists of an aggregation of 44 monolingual and multilingual templated instruction datasets and 19 translated datasets ranging over 114 languages.
-The total size of the Aya dataset is around 513M instances.
-Due to the in-feasibility of fine-tuning our models on the entire dataset, we sampled it down to 2.7M instances, ensuring parity with our sPhinX dataset by selecting equal numbers of samples for each language in our subset.
-Evaluation: We evaluate [^5] Evaluation prompts and other details in Appendix §A.1 our fine-tuned models along with the available Instruction fine-tuned model variants of Mistral-7B and Phi-3-small (IFT [^6] We take the Mistral-7B instruction-tuned variant from here and Phi-3-small variant from here.) on XCOPA Ponti et al.
-(2020), XStoryCloze Lin et al.
-(2022), XWinograd Muennighoff et al.
-(2023)Tikhonov and Ryabinin (2021), Belebele Bandarkar et al.
-(2023), and XQuAD Artetxe et al.
+Base Models: We use Mistral-7B [^3] We specifically use the v1.0 base model from
+https://huggingface.co/mistralai/Mistral-7B-v0.1
+and Phi-3-small Abdin et al. (2024) base model variants and instruction fine-tune them.
+Datasets: Apart from the sPhinX dataset, we use Bactrian Li et al.
+(2023), m-Alpaca Wei et al.
+(2023b) and Aya Singh et al.
+(2024b) instruction datasets for comparative evaluation.
+We also utilize the Azure Translator API [^4] https://azure.microsoft.com/en-us/products/ai-services/ai-translator (sPhinX-T) to translate the original dataset into all our target languages, demonstrating the effectiveness of our Selective Translated Augmentation approach.
+More details about the datasets used for comparative evaluation are present in Appendix §A.2.
+Evaluation: We evaluate [^5] Evaluation prompts and other details in Appendix §A.1 and §A.3.
+our fine-tuned models along with the available base and Instruction fine-tuned model variants of Mistral-7B and Phi-3-small (IFT [^6] We take the Mistral-7B instruction-tuned variant from https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 and Phi-3-small variant from https://huggingface.co/microsoft/Phi-3-small-8k-instruct.) on 4 discriminative tasks; XCOPA Ponti et al.
+(2020)(4-shot), XStoryCloze Lin et al.
+(2022)(4-shot), XWinograd Muennighoff et al.
+(2023)(0-shot), Tikhonov and Ryabinin (2021), Belebele(0-shot) Bandarkar et al.
+(2023), and 2 generative tasks; XQuAD (3-shot) Artetxe et al.
+(2020) and Translation (4-shot) Bojar et al.
+(2014, 2016); Kocmi et al.
 (2023) using the language model evaluation harness Gao et al.
 (2023).
-  • XCOPA: A causal commonsense reasoning dataset in 11 languages, evaluated in a 4-shot prompt setting.
-  • XStoryCloze: A professionally translated version of the English StoryCloze dataset Mostafazadeh et al. (2017) in 10 languages, evaluated in a 4-shot prompt setting.
-  • Belebele: A parallel reading comprehension dataset across 122 languages, with evaluation on a subset of 14 languages in a 0-shot prompt setting.
-  • XQuAD: A QA dataset consisting of professional translations of a subset of SQuAD into 10 languages, evaluated in a 3-shot prompt setting due to context window limitations.
-  • XWinograd: A collection of Winograd Schemas in six languages for cross-lingual commonsense reasoning, evaluated in a 0-shot setting.
-4.2 Fine-Tuning Methodology
-Inspired from  Longpre et al.
-(2023)’s strategies to instruction tune a model, we devise Language-Specific N-shot Guided Instruction fine-tuning (LANG).
-With this approach, we augment a training example by prepending $N$ number of samples of same language as that of the original training example randomly selected from the corresponding few shot set.
-This augmented training example is used for instruction tuning our models.
-Suppose a training example in language $l$ is a pair of $Instruction({I}_{\text{train}}^{l})$ and $Response({R}_{\text{train}}^{l})$.
-We need to prepend $N$ number of shots represented by $({I}_{\text{fewshot}_{1}}^{l},{R}_{\text{fewshot}_{1}}^{l})$, $({I}_{\text{fewshot}_{2}}^{l},{R}_{\text{fewshot}_{2}}^{l})$, …, $({I}_{\text{fewshot}_{N}}^{l},{R}_{\text{fewshot}_{N}}^{l})$.
-For every training example, $N$ is chosen with a probability $p$ as defined in Table 3.
-$\mathcal{T}$ is the instruction tuning templating function which takes $Instruction$ and $Response$ and transforms it to user-assistant format by adding the special tokens.
-Thus, the final training example is:
-$(\bigoplus_{i=1}^{N}\mathcal{T}(I_{\text{fewshot}_{i}}^{l},R_{\text{fewshot}_{i}}^{l})\oplus\mathcal{T}(I_{\text{train}}^{l},R_{\text{train}}^{l}))$
-where $N$ is chosen based on the probability distribution $P(N)$.
-$\bigoplus$ represents combining all few-shot examples, while $\oplus$ indicates concatenating the main training example with the aggregated few-shot examples.
-Once $N$ is selected, an equivalent number of few-shot examples is sampled uniformly at random from the few shot set.
-The maximum context token length for training the models is set at 8192.
-To ensure augmented samples fall within this range, we assign more weights to $N$=0 and $N$=1.
-This strategy is consistently applied when instruction tuning other datasets for both Phi-3-small and Mistral-7B base model.
-The average token length of samples in each datasets as per Phi-3-small model tokenizer is shown in Table 1.
-To assess the effectiveness of each instruction tuning dataset on an equal scale, we conducted a comparative analysis of model performance, fine-tuning each on approximately 8 billion tokens from each dataset using LANG strategy.
-4.3 Hyperparameters and Training Setup
-We used 5 nodes with each node containing 8 A100 GPUs with 80GB VRAM.
-These nodes communicated with each other using InfiniBand  [^7] https://network.nvidia.com/pdf/whitepapers/IB_Intro_WP_190.pdf.
-We use DeepSpeed Rasley et al.
-(2020) to do distributed fine-tuning over these GPUs.
-We use the same hyperparameters (Table 2) to fine-tune both Mistral-7B and Phi-3-small models.
+The number of few-shot selection are inspired from these works Ahuja et al.
+(2023, 2024); Asai et al.
+(2024)
+Apart from generative tasks such as XQuAD, and machine translation, we also evaluate our instruction-tuned models on open-ended generation prompts.
+For this, we use an LLM-based evaluation approach to simulate win rates.
+We use the open-source test set from the Aya Dataset (Singh et al., 2024a), which includes 250 prompts per language across six languages.
+We use GPT-4o as the LLM evaluator to pick the preferred model generation on this test set, and we subsequently compute win rates (%) based on these preferences.
+To avoid a potential bias, we randomize the order of the models during the evaluation.
+The prompt for the evaluator is described in the Appendix: §A.1.
 6 Results
 Figure 3: Performance of Mistral-7B and Phi-3-small when instruction-tuned on 8B tokens across various datasets on different benchmarks.
-Figure 5: Overall performance for both Mistral-7B and Phi-3-small on our dataset and fine-tuning recipe.
-Legend: Base_Model: Base Model, IFT: Instruction Fine-Tuned variant of the Base Model available publicly, Bactrian: Base model trained on Bactrian dataset, mAlpaca: Base Model trained on mAlpaca dataset, Aya: Base Model trained on a subset of the Aya dataset, Sphinx-0s: Base Model trained in a 0-shot fashion on the sPhinX dataset, Sphinx: Base model trained using LANG on sPhinX dataset
-We evaluate the reasoning, question answering, and reading comprehension abilities of the Phi-3-small and Mistral-7B models, instruction-tuned on different multilingual datasets, using various benchmarks and find that fine-tuning on sPhinX provides an improvement of 4.2%pt [^8] A percentage point (pt) is the unit for expressing the absolute difference between two percentage values and 5%pt respectively ( language-wise results are in the Appendix §A.2).
-Additionally as observed in Figure 4, the sPhinX dataset significantly enhances the multilingual performance of the Phi-3-small and Mistral-7B model compared to other datasets even when finetuned on equal number of tokens.
-Furthermore, during instruction tuning on 8B tokens, the models encountered fewer training samples for sPhinX due to its higher average token length per sample, as illustrated in Table 1.
-Due to the code-mixed nature of the instruction along with CoT reasoning explanations, a single sample of sPhinX is notably richer as compared to its counterparts from the other datasets.
-Consequently, even with fewer samples (keeping the number of the tokens same), models trained on sPhinX achieve better performance; thereby demonstrating the per sample efficiency of sPhinX.
-6 Ablation: Improvements from LANG
-To demonstrate the effectiveness of our LANG strategy, we also instruction-tuned the models on sPhinX without any pre-pended shots, referring to this as sPhinX-0s.
-As shown in Figure 5 (with detailed results in Appendix §A.2), models fine-tuned on sPhinX especially Mistral-7B exhibits superior performance compared to its counterparts fine-tuned on other datasets across all benchmarks.
-Moreover, fine-tuning both Phi-3-small and the Mistral-7B on sPhinX using the LANG strategy further boosts the performance by 3.2%pt and 10%pt respectively as compared to the base model fine-tuned without this strategy.
-Surprisingly, even without the LANG strategy, models instruction-tuned on sPhinX still perform better than those tuned on other datasets with LANG.
-Furthermore, employing the LANG strategy leads to additional performance improvements indicating that LANG can effectively enhance the multilingual capabilities of LLMs.
+The solid lines represent the Phi fine-tuned models and the dashed lines represent the Mistral fine-tuned models.
+Model $XC$ $XS$ $XW$ $XQ$ $BL$ $MT^{1}$ $MT^{2}$
+Mistral-7B
+Base Model 0.63 0.68 0.52 0.74 0.24 0.54 0.42
+IFT 0.62 0.73 0.54 0.60 0.47 0.49 0.39
+m-Alpaca 0.55 0.59 0.51 0.46 0.41 0.41 0.39
+Aya 0.68 0.71 0.54 0.66 0.38 0.39 0.37
+Bactrian 0.54 0.67 0.54 0.69 0.26 0.45 0.34
+sPhinX-T 0.61 0.78 0.57 0.78 0.67 0.49 0.38
+sPhinX-0s 0.58 0.58 0.68 0.69 0.67 0.49 0.42
+sPhinX 0.68 0.81 0.71 0.80 0.71 0.55 0.46
+Phi-3-small
+Base Model 0.64 0.78 0.75 0.78 0.65 0.54 0.42
+IFT 0.68 0.79 0.78 0.75 0.70 0.54 0.46
+m-Alpaca 0.68 0.79 0.81 0.77 0.75 0.45 0.39
+Aya 0.65 0.79 0.69 0.83 0.72 0.41 0.40
+Bactrian 0.71 0.82 0.73 0.85 0.77 0.54 0.40
+sPhinX-T 0.70 0.80 0.77 0.78 0.75 0.55 0.44
+sPhinX-0s 0.71 0.81 0.80 0.82 0.79 0.56 0.45
+sPhinX 0.72 0.84 0.87 0.87 0.79 0.56 0.46
+Table 1: Performance of Mistral-7B and Phi-3-small instruction-tuned on various datasets.
+Abbreviations: $XC$ XCOPA (Acc.,4-shot), $XS$ XStoryCloze (Acc.,4-shot) , $XW$ XWinograd (Acc., 0-shot), $XQ$ XQuAD (F1,3-shot), $BL$ Belebele (Acc., 0-shot).
+$MT^{1}$- Translation for x:en (ChrF, 4-shot), $MT^{2}$ Translation for en:x direction (ChrF, 4-shot).
+The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
+We evaluate reasoning, question answering, translation and reading comprehension abilities of the Phi-3-small and Mistral-7B models, instruction-tuned on different multilingual datasets, using various benchmarks and find that fine-tuning on sPhinX provides an average improvement of 39.8% and 11.2% respectively on both the models.
+(Refer to sPhinX-0s in Table 1 for overall results and to Appendix §A.5 for language-wise results).
+Additionally, as observed in Figure 3, the sPhinX dataset significantly enhances the multilingual performance of the Phi-3-small and Mistral-7B model compared to other datasets even when fine-tuned on an equal number of tokens.
+7 Ablations
+7.1 Improvements from LANGIT
+To demonstrate the effectiveness of our LANGIT strategy, we also instruction-tuned the models on sPhinX with 0 shots, referring to this as sPhinX-0s.
+As shown in Table 1 (with detailed results in Appendix §A.5), models fine-tuned on sPhinX especially Mistral-7B exhibit superior performance compared to its counterparts fine-tuned on other datasets across all benchmarks.
+Moreover, fine-tuning both Mistral-7B and the Phi-3-small on sPhinX using the LANGIT strategy further boosts the performance by an average of 15% and 3.2% respectively as compared to the vanilla fine-tuned model (sPhinX -0s) across multilingual benchmarks.
+Furthermore, employing the LANGIT strategy leads to additional performance improvements indicating that LANGIT can effectively enhance the multilingual capabilities of LLMs.
 From the detailed results in Appendix §A.5, we observe no performance regression on high resource languages which normally occurs due to catastrophic forgetting Chang et al.
 (2023).
-We also observe significant performance improvements in medium and low-resource languages such as Arabic, Hindi, Thai, Turkish, Tamil, and Telugu, further showcasing the effectiveness of our dataset and the LANG fine-tuning strategy (Appendix §A.2).
+We also observe significant performance improvements in medium and low-resource languages such as Arabic, Hindi, Thai, Turkish, Tamil, and Telugu, further showcasing the effectiveness of our dataset and the LANGIT fine-tuning strategy (Appendix §A.5).
+7.2 Comparisons with the API translated dataset
+Due to the code-mixed nature of the instruction along with CoT reasoning explanations, a single sample of sPhinX is notably richer as compared to its counterparts from the other datasets.
+This can be observed in the Table 1 for sPhinX-T wherein the sPhinX trained models with the LANGIT strategy outperform the directly translated dataset baselines by an average of 11.7% and 6.3% for both Mistral-7B and Phi-3-small respectively when compared to the sPhinX-T baselines across multilingual benchmarks.
+Consequently, even with fewer samples as compared to the other datasets (keeping the number of the tokens the same), models trained on sPhinX achieve better performance, thereby demonstrating the per-sample efficiency of sPhinX.
+7.3 Simulated Preference Evaluation
+As shown in Table 2, our win-rate experiments reveal that the GPT-4o evaluator predominantly favored outputs generated by the Mistral base model trained on the sPhinX dataset using the LANGIT strategy over other models.
+For the Phi baselines, we observed a higher percentage of TIEs for all the languages except English, where the evaluator rated both outputs equally, rather than favoring a specific model.
+This performance gap between the Mistral and Phi models likely arises from the age of their respective base models.
+Since the Mistral base model is older, it benefits more from additional training on our dataset, whereas the more recently released Phi models are already competitive enough on these benchmarks resulting in preferring both the outputs equally.
+Model ar en po te tu zh
+Mistral-7B
+IFT 75 / 9 / 16 59 / 36 / 5 57 / 27 / 16 62 / 15 / 23 65 / 13 / 22 70 / 26 / 4
+m-Alpaca 85 / 2 / 13 74 / 21 / 5 52 / 33 / 15 69 / 8 / 23 70 / 4 / 25 75 / 15 / 10
+Aya 78 / 11 / 11 85 / 11 / 4 55 / 30 / 15 56 / 18 / 25 62 / 19 / 19 78 / 14 / 8
+Bactrian 82 / 4 / 13 85 / 12 / 3 57 / 31 / 12 56 / 18 / 26 62 / 20 / 18 74 / 14 / 12
+sPhinX-T 61 / 23 / 16 63 / 27 / 10 56 / 24 / 20 56 / 24 / 20 62 / 19 / 19 66 / 23 / 11
+sPhinX-0s 65 / 19 / 16 74 / 20 / 5 55 / 31 / 14 51 / 22 / 27 52 / 36 / 12 68 / 20 / 11
+Phi-3-small
+IFT 46 / 38 / 17 55 / 43 / 2 50 / 44 / 6 39 / 29 / 36 30 / 27 / 43 50 / 44 / 6
+m-Alpaca 46 / 5 / 49 59 / 29 / 12 44 / 21 / 35 33 / 6 / 60 31 / 10 / 59 30 / 6 / 64
+Aya 40 / 18 / 42 80 / 11 / 9 56 / 16 / 28 37 / 18 / 46 25 / 21 / 54 36 / 22 / 41
+Bactrian 32 / 13 / 55 68 / 17 / 15 51 / 14 / 36 24 / 14 / 62 32 / 15 / 53 26 / 11 / 63
+sPhinX-T 35 / 14 / 51 75 / 12 / 13 61 / 12 / 27 23 / 14 / 63 36 / 18 / 47 37 / 12 / 51
+sPhinX-0s 23 / 11 / 66 61 / 17 / 22 32 / 18 / 50 14 / 9 / 77 15 / 10 / 74 14 / 7 / 79
+Table 2: Win rates (%) according to GPT-4o: The first value represents the percentage of outputs where the evaluator preferred the sPhinX and LANGIT trained model.
+The second value indicates the percentage of outputs preferred from the target model.
+The third value reflects cases where the evaluator rated both outputs equally (TIE).
+Benchmarks Base Model sPhinX
+MMLU
+(5-shot)
+0.76 0.75
+HellaSwag
+(5-shot)
+0.81 0.83
+GSM-8k
+(8-shot, CoT)
+0.85 0.77
+MedQA
+(2-shot)
+0.64 0.66
+Arc-C
+(10-shot)
+0.90 0.90
+Arc-E
+(10-shot)
+0.97 0.97
+PIQA
+(5-shot)
+0.84 0.89
+WinoGrande
+(5-shot)
+0.77 0.82
+OpenBookQA
+(10-shot)
+0.86 0.88
+BoolQ
+(2-shot)
+0.82 0.87
+CommonSenseQA
+(10-shot)
+0.80 0.81
+Table 3: Performance of the Phi-3-small base model and the sPhinX tuned model on standard English LLM benchmarks.
 7.4 Regression Analysis on Standard LLM Benchmarks
-It is well-studied that training on multiple languages cause regression on the performance on English language due to catastrophic forgetting Chang et al.
+It is well studied that training in multiple languages causes regression in performance in English due to catastrophic forgetting Chang et al.
 (2023).
-We test this by checking for performance of the Phi-3-small model fine-tuned with sPhinX on English in the multilingual benchmarks we evaluate ((Appendix §A.2) and on popular English-only benchmarks.
-We find that the Phi-3-small sPhinX model maintains its performance in English on the multilingual benchmarks and is also consistently able to maintain performance on standard English benchmarks such as MMLU (5-shot) Hendrycks et al.
+We test this phenomenon for our trained models by checking the performance of the Phi-3-small model fine-tuned with sPhinX on English in the multilingual benchmarks we evaluate ((Appendix §A.5) and on popular English-only benchmarks (Table 3).
+We find that the Phi-3-small fine-tuned on sPhinX maintains its performance in English on the multilingual benchmarks and is also consistently able to maintain performance on standard English benchmarks such as MMLU (5-shot) Hendrycks et al.
 (2021), MedQA (2-shot) Jin et al.
@@
 8 Conclusion
-In this paper, we demonstrated how instruction tuning Phi-3-small and Mistral-7B on sPhinX effectively improve their multilingual capabilities.
-We observed that instruction tuning the models using the sPhinX dataset leads to consistent gains on an average by 4.2%pt for Phi-3-small and 5%pt for Mistral-7B respectively.
-We find that the Phi-3-small model instruction tuned on sPhinX is the best performing model on multilingual benchmarks.
+In this paper, we demonstrated how instruction tuning Mistral-7B and Phi-3-small on sPhinX effectively improve their multilingual capabilities.
+We observed that instruction tuning the models using the sPhinX dataset leads to performance improvement by an average of 39.8% and 11.2% for Mistral-7B and Phi-3-small respectively when compared to their corresponding base models across multilingual benchmarks.
 Moreover, sPhinX exhibits greater sample efficiency and diversity compared to other multilingual instruction tuning datasets.
-Additionally, we proposed a method for further enhancing the model’s performance by utilizing LANG, which supplements the training examples with $N$ samples from a few-shot set, providing the model with additional context to aid in its learning process.
-This further boosts the performance for both Phi-3-small and Mistral-7B by 3.2%pt and 10%pt respectively.
-Models instruction-tuned on sPhinX, exhibit enhanced performance even in languages they have not previously encountered during training.
-Finally, we evaluate the performance of the Phi-3-small model fine-tuned on sPhinX and find that it is able to maintain English performance, suggesting that gains in multilingual performance do not come at the cost of English performance while fine-tuning with sPhinX.
-Through our comprehensive experiments, analyses, and findings, we aim to contribute to the progress of LLMs for multilingual purposes, promoting advancements in natural language processing across a broader spectrum of languages.
+We also proposed LANGIT, a strategy that enhances model performance by incorporating $N$ few-shot examples, boosting results by 15% and 3.2% for Mistral-7B and Phi-3-small, respectively, over vanilla fine-tuning with sPhinX.
+Compared to the sPhinX-T translation baseline, LANGIT yielded gains of 11.7% and 6.3%.
+Models fine-tuned on sPhinX also showed improved performance in unseen languages without degrading English performance.
+We also observed that the GPT-4o win-rate evaluations favored Mistral-7B with LANGIT, while Phi-3-small showed more ties due to its stronger baseline.
+Finally, we plan on releasing a subset of our augmented dataset built on OpenOrca (Open-sPhinX).
 9 Future Work
-We have conducted all experiments using 7B base models in full parameter fine-tuning settings.
-It would be interesting to study the same effect by adaptive learning using LoRA Hu et al.
+All experiments were conducted using 7B base models with full fine-tuning.
+It would be interesting to explore our methods with adaptive fine-tuning techniques like LoRA Hu et al.
 (2022) or PEFT Mangrulkar et al.
-(2022).
-We believe that our strategies could also be effective for models with fewer parameters, leading to a notable improvement in multilingual performance.
-Our LANG strategy involves utilizing $N$ examples from the same language.
-Other ideas that can be explored are incorporating $N$ examples from the same language script to increase the diversity of the sample.
-This approach could be particularly beneficial for enhancing the performance of models in low-resource languages.
+(2022), and on smaller models, where we expect similar gains in multilingual performance.
+Our LANGIT strategy uses $N$ examples from the same language and future work could investigate using $N$ examples from the same script to introduce greater diversity, especially for improving performance in low-resource languages.
+We also observed code-switched data being generated when we employed this strategy to generate data.
+It will be interesting to explore this phenomena in a future study.
 Limitations
@@
 Besides, our study is confined to 7B models, explicitly excluding larger models.
-Despite this limitation, we believe our methodologies are broadly applicable for fine-tuning smaller datasets using techniques like LoRA and PEFT.
-Secondly, our fine-tuning dataset focuses on reasoning tasks and excludes some low-resource languages.
-We evaluated the models’ performance against these reasoning benchmarks.
-However, we did not benchmark our models on generative tasks such as summarization, nor did we evaluate models on hallucination, toxicity, or fairness.
+Despite this limitation, we believe that our methodologies are broadly applicable for fine-tuning smaller datasets using techniques like LoRA and PEFT.
+While some languages were not included, we made a conscious effort to cover a diverse set spanning multiple scripts and language families.
 Ethics Statement
@@
 Hence, it is possible that these issues might propagate to our fine-tuned models as well.
-We are committed to mitigate such risks and strongly advocate for the responsible use of recipes and prevent any unintended negative consequences.
-Acknowledgements
-We thank the Orca team at Microsoft Research for sharing the Orca dataset to implement the synthetic dataset creation recipe.
+We are committed to mitigating such risks and strongly advocate for the responsible use of recipes and prevent any unintended negative consequences.
 Appendix A Appendix
 A.1 Prompt Templates
-Figure 6 is the template for Selective Translation that was used to generate the synthetic data.
+Figure 4 is the template for Selective Translated Augmentation that was used to generate the synthetic data.
 Our reference dataset is in English and the {language} is the target language to generate the data in.
-Figure 7, 8, 9,10 and 11 are the prompts used to evaluate XQuAD, XstoryCloze, Xwinograd, XCOPA and Belebele respectively.
-A.2 Detailed Results
-Tables 4, 5, 6, 7 and 8 show the granular results on our models and dataset.
+Figure 5, 6, 7,8, 9 and 10 are the prompts used to evaluate XQuAD, XstoryCloze, Xwinograd, XCOPA, Belebele and Translation respectively.
+Figure 11 denotes the prompt used for simulating win-rate evaluations.
+A.2 Baseline Datasets For Comparative Evaluation
+  • Bactrian Li et al.
+(2023) is a machine translated dataset of the original alpaca-52k Taori et al.
+(2023) and dolly-15k Conover et al.
+(2023) datasets into 52 languages.
+The instructions for this dataset were translated using a Translation API and then GPT-3.5-Turbo was prompted to generate outputs.
+We fine-tune our models on the complete dataset consisting of 3.4M instances.
+  • m-Alpaca Wei et al.
+(2023b) is a self-instruct dataset that translates seed instructions from English to 11 languages, using GPT-3.5-Turbo for response generation.
+We fine-tune our models on the full dataset, which contains 500k data points.
+  • Aya Singh et al.
+(2024a) contains human-curated prompt-completion pairs in 65 languages, along with 44 monolingual and multilingual instruction datasets and 19 translated datasets across 114 languages, totaling around 513M instances.
+To ensure parity with the sPhinX dataset, we sampled it down to 2.7M instances, ensuring equal representation for each language in our subset.
+A.3 Evaluation Benchmarks
+  • XCOPA: A causal commonsense reasoning dataset in 11 languages, evaluated in a 4-shot prompt setting.
+  • XStoryCloze: A professionally translated version of the English StoryCloze dataset Mostafazadeh et al. (2017) in 10 languages, evaluated in a 4-shot prompt setting.
+  • Belebele: A parallel reading comprehension dataset across 122 languages, with evaluation on a subset of 14 languages in a 0-shot prompt setting.
+  • XQuAD: A QA dataset consisting of professional translations of a subset of SQuAD into 10 languages, evaluated in a 3-shot prompt setting due to context window limitations.
+  • XWinograd: A collection of Winograd Schemas in six languages for cross-lingual commonsense reasoning, evaluated in a 0-shot setting.
+  • Translation: We utilize a subset of WMT14, WMT16 and WMT23 of language pairs (7 languages), with evaluation in a 4-shot setting.
+A.4 Hyperparameters and Training Setup
+We used 5 nodes with each node containing 8 A100 GPUs with 80GB VRAM.
+These nodes communicated with each other using InfiniBand  [^7] https://network.nvidia.com/pdf/whitepapers/IB_Intro_WP_190.pdf.
+We use DeepSpeed Rasley et al.
+(2020) to do distributed fine-tuning over these GPUs.
+We use the same hyperparameters (Table 4) to fine-tune both Mistral-7B and Phi-3-small models.
+A.5 Detailed Results
+Tables 8, 9, 10, 11, 12, 13 and 14 show the granular results on our models and dataset.
 […]
@@
 Figure 9: Belebele evaluation prompt
+[…]
+Translate the following sentence pairs:
+{Source Language}: {Source Phrase} {Target Language}: {Target Phrase}
+Figure 10: Translation evaluation prompt
+[…]
+System: You are a helpful following assistant whose goal is to select the preferred (least wrong) output for a given instruction in {LANGUAGE_NAME}.
+User: Which of the following answers is the best one for given instruction in {LANGUAGE_NAME}. A good answer should follow these rules: 1) It should be in {[LANGUAGE_NAME}.
+2) It should answer the request in the instruction.
+3) It should be factually and semantically comprehensible.
+4) It should be grammatically correct and fluent.
+Instruction: {INSTRUCTION}
+Answer (A): {COMPLETION A}
+Answer (B): {COMPLETION B}
+FIRST provide a one-sentence comparison of the two answers, explaining which you prefer and why.
+SECOND, on a new line, state only ‘Answer (A)’ or ‘Answer (B)’ to indicate your choice.
+If the both answers are equally good or bad, state ‘TIE’.
+Your response should use the format:
+Comparison: <one-sentence comparison and explanation>
+Preferred: <‘Answer (A)’ or ‘Answer (B)’ or ‘TIE’>
+Figure 11: Preference simulation prompt taken from (Üstün et al., 2024) evaluation suite to evaluate our models on free-form generation using GPT-4o.
 Hyperparameter Value
@@
 Context length 8192
-Learning Rate $10*e^{-5}$
+Learning Rate $10^{-5}$
 Scheduler Cosine
@@
 Table 5: Probabilities of selecting number of shots in the LANG strategy
-Language en fr jp pt ru zh avg
-Mistral-7B
-Base Model 0.52 0.47 0.52 0.54 0.54 0.50 0.52
-IFT 0.61 0.57 0.57 0.57 0.60 0.56 0.58
-m-Alpaca 0.61 0.57 0.57 0.57 0.60 0.56 0.58
-Aya 0.55 0.56 0.54 0.54 0.56 0.54 0.55
-Bactrian 0.61 0.57 0.57 0.57 0.60 0.56 0.58
-sPhinX-0s 0.75 0.65 0.68 0.67 0.66 0.65 0.68
-sPhinX 0.80 0.69 0.72 0.70 0.67 0.67 0.71
-Phi-3-small
-Base Model 0.86 0.67 0.73 0.77 0.74 0.72 0.75
-IFT 0.86 0.78 0.72 0.78 0.77 0.75 0.78
-m-Alpaca 0.87 0.76 0.75 0.78 0.76 0.71 0.81
-Aya 0.79 0.61 0.67 0.70 0.70 0.66 0.69
-Bactrian 0.83 0.72 0.71 0.75 0.70 0.68 0.73
-sPhinX-0s 0.89 0.77 0.78 missing0.79 0.81 0.76 0.80
-sPhinX missing0.89 missing0.76 missing0.79 0.78 missing0.82 missing0.77 missing0.84
-Table 4: Language-wise performance of instruction-tuned Mistral-7B and Phi-3-small  models evaluated on XWinograd (0-shot).
-Metric: Accuracy.
-The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
-Language ar de el en es hi ro ru th tr vi zh avg
-Mistral-7B
-Base Model 0.62 0.81 0.64 0.89 0.86 0.65 0.82 0.71 0.59 0.68 0.79 0.76 0.74
-IFT 0.42 0.68 0.33 0.92 0.66 0.5 0.71 0.61 0.38 0.63 0.71 0.68 0.60
-m-Alpaca 0.10 0.75 0.15 0.86 0.82 0.12 0.62 0.68 0.12 0.38 0.52 0.46 0.46
-Aya 0.33 0.73 0.65 0.85 0.8 0.63 0.75 0.67 0.57 0.61 0.75 0.59 0.66
-Bactrian 0.67 0.76 0.26 0.85 0.86 0.74 0.77 0.71 0.59 0.69 0.77 0.65 0.69
-sPhinX-0s 0.54 0.76 0.7 0.88 0.84 0.69 0.77 0.66 0.52 0.64 0.71 0.60 0.69
-sPhinX 0.74 0.87 0.76 0.93 0.90 0.79 0.86 0.77 0.63 0.76 0.88 0.73 0.80
-Phi-3-small
-Base Model 0.68 0.90 0.77 0.93 0.91 0.61 0.84 0.80 0.55 0.73 0.86 0.69 0.78
-IFT 0.71 0.88 0.73 0.92 0.91 0.64 0.84 0.80 0.44 0.70 0.67 0.76 0.75
-m-Alpaca 0.55 0.92 0.74 0.96 0.94 0.68 0.87 0.85 0.50 0.73 0.88 0.66 0.77
-Aya 0.61 0.89 0.84 0.94 0.93 0.80 0.89 0.82 0.73 0.83 0.92 0.79 0.83
-Bactrian 0.81 0.92 0.81 0.95 0.95 0.80 0.91 0.84 0.74 0.82 0.92 0.79 0.85
-sPhinX-0s 0.75 0.89 0.81 0.94 0.94 0.75 0.87 0.79 0.63 0.77 0.88 0.78 0.82
-sPhinX missing0.84 missing0.93 missing0.87 missing0.96 missing0.96 missing0.81 missing0.91 missing0.86 missing0.73 missing0.84 missing0.92 missing0.81 missing0.87
-Table 5: Granular results for XQuAD (3-shot) on our model.
-Metric: F1.
-The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
-Language et ht id it qu sw ta th tr vi zh en avg
-Mistral-7B
-Base Model 0.54 0.51 0.72 0.81 0.49 0.52 0.50 0.53 0.58 0.62 0.78 0.93 0.63
-IFT 0.52 0.52 0.69 0.79 0.50 0.51 0.50 0.54 0.57 0.63 0.75 0.90 0.62
-m-Alpaca 0.51 0.50 0.52 0.63 0.50 0.50 0.50 0.51 0.51 0.49 0.65 0.74 0.55
-Aya 0.57 0.54 0.64 0.67 0.53 0.56 0.57 0.62 0.56 0.61 0.64 0.78 0.61
-Bactrian 0.52 0.50 0.53 0.60 0.49 0.51 0.50 0.51 0.51 0.52 0.52 0.71 0.54
-sPhinX-0s 0.54 0.5 0.58 0.63 0.51 0.55 0.52 0.52 0.54 0.57 0.64 0.8 0.58
-sPhinX missing0.64 0.54 0.73 0.80 0.53 0.61 0.59 0.63 0.67 0.66 0.80 0.91 0.68
-Phi-3-small
-Base Model 0.55 0.51 0.80 0.93 0.52 0.54 0.46 0.56 0.61 0.66 0.86 0.98 0.64
-IFT 0.55 0.57 0.81 0.93 0.53 0.58 0.48 0.60 0.62 0.69 0.88 0.96 0.68
-m-Alpaca 0.53 0.54 0.80 0.92 0.49 0.54 0.51 0.59 0.64 0.68 0.87 0.99 0.68
-Aya 0.60 0.55 0.72 0.83 0.52 0.55 0.52 0.62 0.59 0.69 0.75 0.89 0.65
-Bactrian 0.62 0.56 0.83 0.91 0.54 0.62 0.52 0.66 0.65 0.71 0.86 0.99 0.71
-sPhinX-0s 0.59 0.58 0.85 0.94 0.50 0.60 0.54 0.63 0.69 missing0.72 0.89 0.96 0.71
-sPhinX 0.59 missing0.60 missing0.85 missing0.94 missing0.52 missing0.57 missing0.58 missing0.68 missing0.68 0.71 missing0.90 missing0.98 missing0.72
-Table 6: Granular results for XCOPA (4-shot) on our model.
-Metric: Accuracy.
-The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
-Language ar en es eu hi id my ru sw te zh avg
-Mistral-7B
-Base Model 0.65 0.89 0.83 0.56 0.62 0.76 0.52 0.81 0.56 0.52 0.80 0.68
-IFT 0.70 0.96 0.92 0.54 0.69 0.79 0.57 0.90 0.58 0.54 0.88 0.73
-m-Alpaca 0.53 0.73 0.7 0.51 0.51 0.57 0.50 0.66 0.52 0.52 0.71 0.59
-Aya 0.64 0.86 0.81 0.56 0.71 0.73 0.60 0.82 0.67 0.60 0.81 0.71
-Bactrian 0.69 0.82 0.74 0.52 0.59 0.76 0.54 0.73 0.62 0.61 0.76 0.67
-sPhinX-0s 0.57 0.66 0.64 0.47 0.56 0.61 0.50 0.62 0.56 0.52 0.69 0.58
-sPhinX 0.83 0.96 0.94 0.57 missing0.84 0.87 missing0.61 0.91 missing0.80 missing0.69 0.94 0.81
-Phi-3-small
-Base Model 0.80 0.98 0.96 0.61 0.72 0.92 0.53 0.96 0.61 0.55 0.94 0.78
-IFT 0.81 0.98 0.96 0.61 0.75 0.92 0.56 0.96 0.61 0.53 0.94 0.79
-m-Alpaca 0.81 0.99 0.98 0.58 0.76 0.93 0.52 0.97 0.64 0.54 0.96 0.79
-Aya 0.77 0.98 0.97 0.57 0.77 0.93 0.53 0.96 0.74 0.56 0.94 0.79
-Bactrian 0.83 0.98 0.98 0.61 0.83 0.94 0.54 0.97 0.79 0.63 0.94 0.82
-sPhinX-0s 0.84 0.98 0.97 missing0.64 0.77 0.95 0.52 0.96 0.74 0.57 0.95 0.81
-sPhinX missing0.86 missing0.99 missing0.99 0.61 0.82 missing0.96 0.54 missing0.98 0.74 0.61 missing0.97 missing0.82
-Table 7: Granular results for XStoryCloze (4-shot) on our model.
-Metric: Accuracy.
-The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
-Language ar de es en fi fr hi it jp ko ta te vi zh avg
-Mistral-7B
-Base Model 0.25 0.23 0.23 0.24 0.23 0.23 0.26 0.24 0.26 0.23 0.23 0.25 0.26 0.25 0.24
-IFT 0.32 0.60 0.62 0.74 0.36 0.62 0.32 0.61 0.43 0.47 0.27 0.27 0.39 0.58 0.47
-m-Alpaca 0.32 0.50 0.53 0.56 0.45 0.51 0.27 0.51 0.40 0.41 0.26 0.26 0.33 0.48 0.41
-Aya 0.34 0.43 0.43 0.48 0.38 0.47 0.35 0.44 0.4 0.36 0.27 0.25 0.37 0.42 0.38
-Bactrian 0.24 0.27 0.25 0.25 0.26 0.27 0.24 0.28 0.26 0.26 0.23 0.23 0.34 0.28 0.26
-sPhinX-0s 0.64 0.75 0.75 0.82 0.66 0.79 0.53 0.73 0.69 0.66 0.48 0.44 0.66 0.75 0.67
-sPhinX 0.69 0.80 0.69 0.87 0.71 0.82 missing0.60 0.79 0.73 0.73 missing0.56 missing0.48 0.70 0.80 0.71
-Phi-3-small
-Base Model 0.54 0.87 0.85 0.92 0.58 0.86 0.41 0.86 0.70 0.58 0.26 0.30 0.62 0.82 0.65
-IFT 0.63 0.89 0.88 0.93 0.63 0.88 0.48 0.88 0.77 0.68 0.32 0.32 0.68 0.85 0.70
-m-Alpaca 0.65 0.92 0.90 0.94 0.74 0.91 0.54 0.90 0.80 0.70 0.47 0.45 0.72 0.84 0.75
-Aya 0.58 0.86 0.85 0.91 0.65 0.87 0.50 0.86 0.76 0.67 0.37 0.35 0.69 0.84 0.70
-Bactrian 0.67 0.88 0.88 0.92 0.70 0.88 0.51 0.86 0.77 0.70 0.37 0.37 0.74 0.86 0.72
-sPhinX-0s 0.73 0.91 0.90 0.93 0.75 0.92 0.57 0.91 0.82 missing0.82 0.45 0.40 0.76 0.89 0.77
-sPhinX missing0.74 missing0.93 missing0.91 missing0.94 missing0.77 missing0.93 0.58 missing0.92 missing0.84 0.76 missing0.46 0.40 missing0.78 missing0.89 missing0.79
-Table 8: Granular results for Belebele (0-shot) on our model.
-Metric: Accuracy.
-The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
-Model XC XS XW XQ BL
-Mistral-7B
-Base Model 0.63 0.68 0.52 0.74 0.24
-IFT 0.62 0.73 0.54 0.60 0.47
-m-Alpaca 0.55 0.59 0.51 0.46 0.41
-Aya 0.68 0.71 0.54 0.66 0.38
-Bactrian 0.54 0.67 0.54 0.69 0.26
-sPhinX-0s 0.58 0.58 0.68 0.69 0.67
-sPhinX 0.68 0.81 0.71 0.80 0.71
-Phi-3-small
-Base Model 0.64 0.78 0.75 0.78 0.65
-IFT 0.68 0.79 0.78 0.75 0.70
-m-Alpaca 0.68 0.79 0.81 0.77 0.75
-Aya 0.65 0.79 0.69 0.83 0.72
-Bactrian 0.71 0.82 0.73 0.85 0.77
-sPhinX-0s 0.71 0.81 0.80 0.82 0.79
-sPhinX 0.72 0.84 0.87 0.87 0.79
-Table 9: Performance of Mistral-7B and Phi-3-small instruction-tuned for 10000 training steps on various datasets.
-Abbreviations: XC - XCOPA, XS - XStoryCloze, XW - XWinograd, XQ - XQuAD, BL - Belebele.
-The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
-Benchmarks Base Model sPhinX
-MMLU
-(5-shot)
-0.76 0.75
-HellaSwag
-(5-shot)
-0.81 0.83
-GSM-8k
-(8-shot, CoT)
-0.85 0.77
-MedQA
-(2-shot)
-0.64 0.66
-Arc-C
-(10-shot)
-0.90 0.90
-Arc-E
-(10-shot)
-0.97 0.97
-PIQA
-(5-shot)
-0.84 0.89
-WinoGrande
-(5-shot)
-0.77 0.82
-OpenBookQA
-(10-shot)
-0.86 0.88
-BoolQ
-(2-shot)
-0.82 0.87
-CommonSenseQA
-(10-shot)
-0.80 0.81
-Table 10: Performance of the Phi-3-small base model and the sPhinX tuned model on standard English LLM benchmarks.
+Dataset Average Token
+Length/Sample
+Aya 2240
+Bactrian 2465
+m-Alpaca 1620
+sPhinX-0s 544
+sPhinX 3100
+Table 6: Average Token Length in each dataset
 High-Resource (100k)
@@
 Table 7: Language distribution and samples across three tiers
+Figure 12: Some examples of input queries and its counterpart existing in the hindi version of the Multialpaca dataset and if it was generated using the Selective Translated Augmentation strategy.
+Again, we observe that the samples generated using Selective Translated Augmentation translate only the required amount of information as controlled via prompting whereas in Multialpaca the translations are direct translations where only a part of the instructions have been followed to translate the input queries.
+Language en fr jp pt ru zh avg
+Mistral-7B
+Base Model 0.52 0.47 0.52 0.54 0.54 0.50 0.52
+IFT 0.61 0.57 0.57 0.57 0.60 0.56 0.58
+m-Alpaca 0.61 0.57 0.57 0.57 0.60 0.56 0.58
+Aya 0.55 0.56 0.54 0.54 0.56 0.54 0.55
+Bactrian 0.61 0.57 0.57 0.57 0.60 0.56 0.58
+sPhinX-T 0.58 0.61 0.57 0.53 0.54 0.57 0.57
+sPhinX-0s 0.75 0.65 0.68 0.67 0.66 0.65 0.68
+sPhinX 0.80 0.69 0.72 0.70 0.67 0.67 0.71
+Phi-3-small
+Base Model 0.86 0.67 0.73 0.77 0.74 0.72 0.75
+IFT 0.86 0.78 0.72 0.78 0.77 0.75 0.78
+m-Alpaca 0.87 0.76 0.75 0.78 0.76 0.71 0.81
+Aya 0.79 0.61 0.67 0.70 0.70 0.66 0.69
+Bactrian 0.83 0.72 0.71 0.75 0.70 0.68 0.73
+sPhinX-T 0.87 0.74 0.74 0.77 0.77 0.72 0.77
+sPhinX-0s 0.88 0.75 0.78 missing0.79 0.81 0.76 0.80
+sPhinX missing0.89 missing0.76 missing0.79 missing0.79 missing0.82 missing0.77 missing0.84
+Table 8: Language-wise performance of instruction-tuned Mistral-7B and Phi-3-small  models evaluated on XWinograd (0-shot).
+Metric: Accuracy.
+The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
+Language ar de el en es hi ro ru th tr vi zh avg
+Mistral-7B
+Base Model 0.62 0.81 0.64 0.89 0.86 0.65 0.82 0.71 0.59 0.68 0.79 0.72 0.73
+IFT 0.42 0.68 0.33 0.92 0.66 0.5 0.71 0.61 0.38 0.63 0.71 0.68 0.60
+m-Alpaca 0.10 0.75 0.15 0.86 0.82 0.12 0.62 0.68 0.12 0.38 0.52 0.46 0.46
+Aya 0.33 0.73 0.65 0.85 0.80 0.63 0.75 0.67 0.57 0.61 0.75 0.59 0.66
+Bactrian 0.67 0.76 0.26 0.85 0.86 0.74 0.77 0.71 0.59 0.69 0.77 0.65 0.69
+sPhinX-T 0.71 0.83 0.75 0.92 0.89 0.77 0.84 0.75 0.60 0.77 0.86 0.63 0.78
+sPhinX-0s 0.54 0.76 0.70 0.88 0.84 0.69 0.77 0.66 0.52 0.64 0.71 0.60 0.69
+sPhinX 0.74 0.87 0.77 0.93 0.90 0.79 0.86 0.77 0.63 0.77 0.88 0.73 0.80
+Phi-3-small
+Base Model 0.68 0.90 0.77 0.93 0.91 0.61 0.84 0.80 0.55 0.73 0.86 0.69 0.78
+IFT 0.71 0.88 0.73 0.92 0.91 0.64 0.84 0.80 0.44 0.70 0.67 0.76 0.75
+m-Alpaca 0.55 0.92 0.74 0.96 0.94 0.68 0.87 0.85 0.50 0.73 0.88 0.66 0.77
+Aya 0.61 0.89 0.84 0.94 0.93 0.80 0.89 0.82 0.73 0.83 0.91 0.79 0.83
+Bactrian 0.81 0.92 0.81 0.95 0.95 0.80 0.90 0.84 0.72 0.82 0.91 0.79 0.85
+sPhinX-T 0.80 0.91 0.82 0.95 0.94 0.80 0.90 0.83 0.69 0.81 0.91 0.73 0.78
+sPhinX-0s 0.75 0.89 0.81 0.94 0.94 0.75 0.87 0.79 0.63 0.77 0.88 0.78 0.82
+sPhinX missing0.84 missing0.93 missing0.87 missing0.96 missing0.96 missing0.81 missing0.91 missing0.86 missing0.73 missing0.84 missing0.92 missing0.81 missing0.87
+Table 9: Granular results for XQuAD (3-shot) on our model.
+Metric: F1.
+The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
+Language et ht id it qu sw ta th tr vi zh en avg
+Mistral-7B
+Base Model 0.54 0.51 0.72 0.81 0.49 0.52 0.50 0.53 0.58 0.62 0.78 0.93 0.63
+IFT 0.52 0.52 0.69 0.79 0.50 0.51 0.50 0.54 0.57 0.63 0.75 0.90 0.62
+m-Alpaca 0.51 0.50 0.52 0.63 0.50 0.50 0.50 0.51 0.51 0.49 0.65 0.74 0.55
+Aya 0.57 0.54 0.64 0.67 0.53 0.56 0.57 0.62 0.56 0.61 0.64 0.78 0.61
+Bactrian 0.52 0.50 0.53 0.60 0.49 0.51 0.50 0.51 0.51 0.52 0.52 0.71 0.54
+sPhinX-T 0.57 0.50 0.64 0.72 0.50 0.50 0.58 0.57 0.57 0.62 0.71 0.83 0.61
+sPhinX-0s 0.54 0.5 0.58 0.63 0.51 0.55 0.52 0.52 0.54 0.57 0.64 0.8 0.58
+sPhinX missing0.64 0.54 0.73 0.80 0.53 missing0.61 0.59 0.63 0.67 0.66 0.80 0.91 0.68
+Phi-3-small
+Base Model 0.55 0.51 0.80 0.93 0.52 0.54 0.46 0.56 0.61 0.66 0.86 0.98 0.64
+IFT 0.55 0.57 0.81 0.93 0.53 0.58 0.48 0.60 0.62 0.69 0.88 0.96 0.68
+m-Alpaca 0.53 0.54 0.80 0.92 0.49 0.54 0.51 0.59 0.64 0.68 0.87 0.99 0.68
+Aya 0.60 0.55 0.72 0.83 0.52 0.55 0.52 0.62 0.59 0.69 0.75 0.89 0.65
+Bactrian 0.62 0.56 0.83 0.91 0.52 0.60 0.52 0.66 0.65 0.71 0.86 0.98 0.70
+sPhinX-T 0.55 0.58 0.83 0.93 0.51 0.57 0.56 0.66 0.68 0.68 0.87 0.98 0.70
+sPhinX-0s 0.59 0.58 0.84 0.93 0.50 0.60 0.54 0.63 0.68 missing0.72 0.89 0.96 0.71
+sPhinX 0.59 missing0.60 missing0.85 missing0.94 missing0.52 0.57 missing0.58 missing0.68 missing0.69 0.71 missing0.90 missing0.99 missing0.72
+Table 10: Granular results for XCOPA (4-shot) on our model.
+Metric: Accuracy.
+The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
+Language ar en es eu hi id my ru sw te zh avg
+Mistral-7B
+Base Model 0.65 0.89 0.83 0.56 0.62 0.76 0.52 0.81 0.56 0.52 0.80 0.68
+IFT 0.70 0.95 0.92 0.54 0.69 0.79 0.57 0.90 0.58 0.54 0.88 0.73
+m-Alpaca 0.53 0.73 0.70 0.51 0.51 0.57 0.50 0.66 0.52 0.52 0.71 0.59
+Aya 0.64 0.86 0.81 0.56 0.71 0.73 0.60 0.82 0.67 0.60 0.81 0.71
+Bactrian 0.69 0.82 0.74 0.52 0.59 0.76 0.54 0.73 0.62 0.61 0.76 0.67
+sPhinX-T 0.78 0.92 0.87 0.56 0.80 0.81 0.66 0.86 0.74 0.70 0.86 0.78
+sPhinX-0s 0.57 0.66 0.64 0.47 0.56 0.61 0.50 0.62 0.56 0.52 0.69 0.58
+sPhinX 0.83 0.96 0.94 0.57 missing0.84 0.87 missing0.67 0.91 missing0.80 missing0.69 0.94 0.81
+Phi-3-small
+Base Model 0.80 0.98 0.96 0.61 0.72 0.92 0.53 0.96 0.61 0.55 0.94 0.78
+IFT 0.81 0.98 0.96 0.61 0.75 0.92 0.56 0.96 0.61 0.53 0.94 0.79
+m-Alpaca 0.81 0.98 0.98 0.58 0.76 0.93 0.52 0.97 0.64 0.54 0.96 0.79
+Aya 0.77 0.98 0.97 0.57 0.77 0.93 0.53 0.96 0.74 0.56 0.94 0.79
+Bactrian 0.83 0.98 0.98 0.61 0.83 0.94 0.54 0.97 0.79 0.63 0.94 0.82
+sPhinX-T 0.84 0.98 0.98 0.60 0.80 0.95 0.52 0.96 0.72 0.60 0.85 0.80
+sPhinX-0s 0.84 0.98 0.97 missing0.64 0.77 0.95 0.52 0.96 0.74 0.57 0.95 0.81
+sPhinX missing0.86 missing0.99 missing0.99 0.61 0.82 missing0.96 0.54 missing0.98 0.74 0.61 missing0.97 missing0.82
+Table 11: Granular results for XStoryCloze (4-shot) on our model.
+Metric: Accuracy.
+The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
+Language ar de es en fi fr hi it jp ko ta te vi zh avg
+Mistral-7B
+Base Model 0.25 0.23 0.23 0.24 0.23 0.23 0.26 0.24 0.26 0.23 0.23 0.25 0.26 0.25 0.24
+IFT 0.32 0.60 0.62 0.74 0.36 0.62 0.32 0.61 0.43 0.47 0.27 0.27 0.39 0.58 0.47
+m-Alpaca 0.32 0.50 0.53 0.56 0.45 0.51 0.27 0.51 0.40 0.41 0.26 0.26 0.33 0.48 0.41
+Aya 0.34 0.43 0.43 0.48 0.38 0.47 0.35 0.44 0.4 0.36 0.27 0.25 0.37 0.42 0.38
+Bactrian 0.24 0.27 0.25 0.25 0.26 0.27 0.24 0.28 0.26 0.26 0.23 0.23 0.34 0.28 0.26
+sPhinX-T 0.66 0.74 0.71 0.81 0.66 0.76 0.57 0.73 0.66 0.68 0.54 0.46 0.66 0.71 0.68
+sPhinX-0s 0.64 0.75 0.75 0.82 0.66 0.79 0.53 0.73 0.69 0.66 0.48 0.44 0.66 0.75 0.67
+sPhinX 0.69 0.80 0.69 0.87 0.71 0.82 missing0.60 0.79 0.73 0.73 missing0.56 missing0.48 0.70 0.80 0.71
+Phi-3-small
+Base Model 0.54 0.87 0.85 0.92 0.58 0.86 0.41 0.86 0.70 0.58 0.26 0.30 0.62 0.82 0.65
+IFT 0.63 0.89 0.88 0.93 0.63 0.88 0.48 0.88 0.77 0.68 0.32 0.32 0.68 0.85 0.70
+m-Alpaca 0.65 0.92 0.90 0.94 0.74 0.91 0.54 0.90 0.80 0.70 0.47 0.45 0.72 0.84 0.75
+Aya 0.58 0.86 0.85 0.91 0.65 0.87 0.50 0.86 0.76 0.67 0.37 0.35 0.69 0.84 0.70
+Bactrian 0.67 0.88 0.88 0.92 0.70 0.88 0.51 0.86 0.77 0.70 0.37 0.37 0.74 0.86 0.72
+sPhinX-T 0.71 0.90 0.90 0.93 0.73 0.90 0.56 0.89 0.82 0.75 0.42 0.38 0.75 0.87 0.75
+sPhinX-0s 0.73 0.91 0.90 0.93 0.75 0.92 0.57 0.91 0.82 missing0.82 0.45 0.40 0.76 0.89 0.77
+sPhinX missing0.74 missing0.93 missing0.91 missing0.94 missing0.77 missing0.93 0.58 missing0.92 missing0.84 0.76 missing0.46 0.40 missing0.78 missing0.89 missing0.79
+Table 12: Granular results for Belebele (0-shot) on our model.
+Metric: Accuracy.
+The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
+Language ar fr de ro ja ru zh avg
+Mistral-7B
+Base Model 0.48 0.63 0.65 0.56 0.43 0.54 0.48 0.54
+IFT 0.37 0.61 0.61 0.58 0.28 0.5 0.49 0.49
+m-Alpaca 0.34 0.51 0.51 0.46 0.29 0.36 0.41 0.41
+Aya 0.30 0.50 0.51 0.46 0.24 0.35 0.41 0.39
+Bactrian 0.40 0.55 0.52 0.51 0.34 0.44 0.42 0.45
+sPhinX-T 0.39 0.60 0.60 0.54 0.39 0.49 0.48 0.49
+sPhinX-0s 0.40 0.57 0.61 0.53 0.40 0.51 0.44 0.49
+sPhinX 0.45 0.63 0.64 0.59 0.45 0.55 0.52 0.54
+Phi-3-small
+Base Model 0.48 0.61 0.65 0.57 0.45 0.52 0.53 0.54
+IFT 0.43 0.62 0.63 0.52 0.41 0.49 0.50 0.54
+m-Alpaca 0.44 0.60 0.61 0.56 0.32 0.44 0.19 0.45
+Aya 0.44 0.56 0.57 0.54 0.12 0.45 0.17 0.41
+Bactrian 0.48 0.63 0.65 0.59 0.45 0.52 0.52 0.54
+sPhinX-T 0.48 0.64 0.65 0.59 0.46 0.53 0.52 0.55
+sPhinX-0s 0.49 0.63 0.65 0.59 0.46 0.54 0.53 0.56
+sPhinX 0.49 0.64 0.66 0.60 0.46 0.54 0.53 0.56
+Table 13: Granular results for Translation for language to English direction (4-shot) on our model.
+Metric: ChrF.
+The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
+Language ar fr de ro ja ru zh avg
+Mistral-7B
+Base Model 0.29 0.60 0.54 0.52 0.21 0.34 0.47 0.42
+IFT 0.16 0.58 0.57 0.48 0.17 0.29 0.43 0.38
+m-Alpaca 0.15 0.62 0.66 0.40 0.12 0.48 0.44 0.41
+Aya 0.12 0.54 0.63 0.48 0.16 0.39 0.42 0.39
+Bactrian 0.14 0.51 0.49 0.44 0.10 0.35 0.40 0.38
+sPhinX-T 0.31 0.58 0.53 0.47 0.20 0.30 0.26 0.38
+sPhinX-0s 0.30 0.60 0.55 0.51 0.22 0.31 0.45 0.42
+sPhinX 0.35 0.61 0.60 0.55 0.26 0.36 0.49 0.46
+Phi-3-small
+Base Model 0.31 0.63 0.60 0.43 0.24 0.30 0.45 0.42
+IFT 0.29 0.61 0.58 0.39 0.21 0.27 0.41 0.46
+m-Alpaca 0.39 0.60 0.58 0.31 0.11 0.24 0.47 0.38
+Aya 0.17 0.62 0.56 0.45 0.22 0.28 0.46 0.39
+Bactrian 0.30 0.60 0.56 0.47 0.18 0.24 0.44 0.40
+sPhinX-T 0.33 0.63 0.60 0.48 0.24 0.31 0.48 0.43
+sPhinX-0s 0.32 0.63 0.61 0.50 0.27 0.36 0.49 0.45
+sPhinX 0.35 0.64 0.61 0.51 0.26 0.33 0.49 0.46
+Table 14: Granular results for Translation for English to language direction (4-shot) on our model.
+Metric: ChrF.
+The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
 Code Languages Script Data